A Complete Bibliography of Philosophy of Science: 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 November 2014 Version 1.01

Title word cross-reference

1500 [530].

A. [253, 380]. Aaron [392]. Abduction [352]. Absolute [447]. Absorption [362]. Acquisition [403]. Action [356]. Activity [112]. Actual [405, 310]. Adams [51]. Ado [396]. Advanced [138]. Advancement [457]. Advances [90]. after [251]. against [47]. Albert [177]. Alfred [372, 135]. Amateurs [525]. Ambiguities [115]. America [51]. American [503]. Americans [51]. Analysed [42, 497]. Analysis [180, 241, 484, 477, 476, 414, 242, 523, 330, 243, 12, 266, 27, 305, 514, 483]. Anatomy [341]. Anderson [481]. Animadversion [220]. Annotationes [81]. Announcements [108, 129]. Anthropological [412]. Anthropology [473]. Anthropomorphic [101]. Anton [454]. Appeal [306, 371]. Application [34]. Applications [90]. Applied [503, 504, 501, 507, 343]. Approach [321]. Art [390, 125]. Arthur [498, 369, 373, 374]. Aspects [449, 377, 405, 441]. Astronomy [81]. Atemporal [421]. Atom [52]. Atomic [316]. Attempt [374]. Attitude [222, 513]. Attributes [206]. Attwood [497]. Auguste [66]. Authoritarian [374].

B [51, 52, 122]. Back [166, 181, 194, 204, 214, 227, 239, 249, 326, 338, 346, 359, 467]. Background [295]. Bacon [481]. Baker [385]. Basic [193, 504, 230, 361, 184]. Basis [163, 128, 60, 224]. Baudouin [65]. Bayle [147]. Beauty [253]. Becoming [231]. Behavior [247, 78, 208, 60, 304]. Behaviorist [330]. Behavioristic [459]. Being [158, 231]. Bell [136]. Bergmann [479]. Bertrand [106, 305, 322]. between [151, 6]. Biological [278, 365, 80, 163, 43, 407, 127, 87, 234, 13]. Biology [424, 161, 441, 90]. Biotypology [224]. Birkhoff [290]. Blamey [15]. Bliven [374]. Bond [89]. Book [526, 511, 453, 454, 482, 499, 513, 514, 466, 498, 512, 105, 483, 484, 325, 528, 530, 510, 65, 66, 304, 481, 15, 16, 17, 18, 19, 20, 21, 22, 35, 36, 37, 38, 49, 50, 51, 52, 53, 67, 68, 82, 83, 84, 85, 86, 107, 123, 124, 125, 126, 127, 128, 176, 177, 306, 307, 368, 369, 370, 371, 372, 373, 374, 383, 384, 385, 386, 387, 388, 389, 390, 391, 392, 393, 394, 395, 89, 527, 189, 190, 90, 91, 92, 93, 106, 134, 135, 136, 137, 138, 148, 497, 525, 452, 122, 81, 305, 48]. Books [149, 164, 175, 191, 226]. Boole [484]. Bowdery [354]. Breakdown [406]. Bridgman [189]. Bright [37]. Bruce [374]. Bryson [51, 384]. Buchler [22]. Burtt [369].

C [453, 513, 68, 389, 134, 148, 497, 48]. Calculus [180]. Calkins [127]. Called [343]. Can [462, 461]. Canons [27]. Capture [374]. Carl [125]. Carnap [323, 357, 523, 190]. Carroll [48]. Case [415]. Cassirer [233]. Catalysis [234]. Categorial [414]. Category [405]. Causal [300, 169]. Causality [143]. Causation [398, 43, 170, 30, 234]. Cells [126]. Centering [355]. Certain [237]. Certainty [59, 4]. Cézanne [188]. Challenge [303]. Chambers [68]. Charge [132]. Charges [132]. Charles [325, 65, 8, 22]. Chemical [89]. Chemistry [75, 299]. Chez [66]. Christian [527]. Churchman [494]. Clarke [51]. Class [180, 262, 220, 293, 118]. Clinical [473, 507]. Coherence [480, 197]. Collingwood [370]. Commands [312]. Comments [323, 495, 464, 463, 243, 460]. Common [258, 157]. Communication [382, 509]. Comparative [341]. Comparison [103]. Component [157]. Compound [324]. Comte [66]. Concept [193, 353, 102, 380, 523, 425, 152]. Conception [211]. Concepts [297, 518, 458]. Conceptual [302, 282]. Concerning [436, 192, 293, 302]. Condition [436]. Confirmation [289]. Conflicts [406]. Conn [52]. Consequences [11]. Conservation [115]. Contemporary [332, 72]. Content [242, 243]. Contexts [465]. Continuum [334]. Contradictory [33]. Contribution [439]. Control [50]. Conventionalism [198]. Conventions [381, 141]. Cooperative [436]. Corism [343]. Cosmic [206]. Cosmical [178]. Cosmo [53]. Cosmo-Retardation [53]. Cosmological [147]. Cosmology [81]. Cousins [383]. Cow [17]. Cratylus [252]. Creative [209]. Creators [121]. Crisis [307]. Criteria [478]. Critical [27]. Criticism [310]. Critique [18]. Crossroads [395]. Crowther [123]. Crystals [52]. Cultural [274, 356]. Current [174]. Cybernetics [499]. Cyclotron [52].

D [498, 67, 176, 306, 371]. D. [17]. Dagobert [176]. Data [381, 183, 256, 378]. David [464]. Deaf [509]. Decades [363]. Decency [225]. Decisions [230]. Decline [56]. Decrassifying [99]. Deductive [135]. Defence [310]. Defense [71]. Definition [422, 289]. Definitions [337]. Degree [289]. Deming [50]. Democratic [472]. Departure [426]. Descent [269]. Descriptions [465]. Descriptive [292]. Destiny [183]. Determination [362]. Development [448, 136]. Dewey [374, 99]. Dialectic [363]. Dialectical [489, 281]. Dictionary [176, 68]. Different [61]. Differential [137]. Difficulties [302, 282]. Diffusion [87]. Dimensions [423]. Directive [365]. Directiveness [407]. Discussion [440, 362]. Distinction [6]. Do [438]. Doctor [255]. Domain [368]. Dorothy [525]. Doubt [447]. Doubts [400]. Dr. [418]. Dualism [361]. Ducasse [66, 148]. Dynamics [88].

East [521]. East-West [521]. Economic [448, 304, 302, 12, 282]. Economics [449]. Eddington [498]. Editorial [94, 96, 251, 431, 471]. Edman [374]. Education [374]. Educational [496, 34]. Edwin [369]. Effect [425]. Einstein [290]. Eisenhart [137]. Electrical [353]. Electron [52]. Electronic [132]. Elementary [138]. Elements [292]. Elimination [244]. Emergence [272, 75]. Emotions [330, 77]. Emphasis [222]. Empirical [480, 286, 62, 222, 372]. Empiricism [207, 192, 120, 400, 139, 8, 22]. Empiricist [162, 116]. Empiricus [120]. Enemies [452]. Energy [193, 115, 131]. Engineering [505]. Enquiry [341]. Entities [310]. Entity [377, 405]. Epistemic [245]. Epistemological [419, 363]. Epistemology [192, 399]. Equal [114]. Erhardt [81, 81]. Erhardt-Siebold [81]. Erigena [81]. Ernest [511]. Erratum [165]. Erwin [393]. ESP [236]. Essai [66]. Essay [84, 252]. Essays [510, 389]. Essential [207, 436]. Esthetic [301]. Eternal [345]. Ethic [367]. Ethics [408, 163, 62, 415, 526, 325]. Ethnological [76]. Euclid [333]. Euclideans [333]. Evening [492]. Events [362, 512]. Evolution [80, 425, 299, 13, 128]. Excluded [74]. Exclusion [271]. Existence [381, 316]. Existential [245]. Experience [287, 335, 306, 371]. Experiment [132]. Experimental [296, 413, 124]. Experiments [76]. Expert [472]. Explanation [140, 442, 463]. Exploitation [121]. Extensive [334]. External [186, 29].

F [453, 16, 68, 84, 127, 497]. Fact [223]. Factor [246, 267]. Factors [300]. Facts [237, 61]. Faith [528]. Fallacy [522, 10]. Feigl [514]. Felix [138]. Feynman [362]. Field [353]. Fields [286]. Figure [488]. Fluxive [10]. Footnote [412]. Force [165, 142, 153]. Forces [151]. Foreword [247].

Formal [474]. Formalism [298]. Formation [501]. Foundations [366, 335]. Fourth [488]. Fowler [91]. Framework [62]. Francis [481]. Frank [482]. Freedom [510, 40, 156, 384]. Front [23, 39, 54, 70, 95, 109, 130, 150, 167, 182, 195, 205, 215, 228, 240, 250, 257, 264, 270, 277, 283, 294, 308, 315, 327, 339, 347, 360, 375, 397, 410, 417, 429, 444, 456, 470, 485, 500, 516, 532]. Fulton [481]. Function [291]. Functional [171, 405]. Functions [202]. Further [293]. Future [362, 117, 280].

G [49, 52, 123, 127, 177, 370, 92]. G. [211]. Games [304]. Gamow [49]. General [508, 446]. Generalization [378]. Genes [134]. Geometry [165, 88, 198, 390, 142, 153, 137]. George [484, 354]. Gerard [126]. Gerber [368]. Gerth [389]. Gestaltism [508]. Getting [51]. Ghyka [390]. Gnosology [450]. Goldschmidt [128]. Good [388]. Goodness [253]. Gotshalk [67]. Grading [260]. Greatest [157]. Greek [82]. Group [483, 483]. Guggenheim [91]. Gustaf [85].

H [513, 530, 481, 17, 389, 211, 91, 134]. Habits [187]. Hans [387]. Harold [38]. Harry [526]. Having [97]. Health [51]. Hegel [18]. Heisenberg [111]. Henry [388, 395, 527]. Herbert [514, 466]. Heuristic [321]. Historical [256, 329]. Historicism [272]. Historico [145]. Historico-Sociological [145]. History [80, 287, 64, 525, 415, 364, 13, 86, 124]. Holism [272]. Hollingworth [526]. Hook [374]. Horace [510]. Hugh [86]. Hughes [68]. Human [71, 388, 60]. Humanism [285, 521, 83]. Humanities [511, 453]. Hume [487]. Hyam [392]. Hypothesis [301, 146, 311].

Idea [370]. Ideal [478]. Idealism [427, 199, 98, 210, 160]. Identity [232, 263, 223]. Ideological [292]. Ideology [490]. Ignorance [114]. II [296, 87]. III [297]. Illustrated [252]. Imagination [209]. Immediate [306, 371]. Imperatives [254, 319]. Implication [221]. Importance [271]. Incompatibilities [406]. Inconceivable [458]. Incorrigibility [184]. Indeterminacy [448]. Indeterminism [209]. Indexical [202]. Induction [323, 451, 495, 494, 7, 269, 186]. Inductive [288, 357]. Inference [493, 169]. Infinite [269]. Information [69, 154, 213, 229, 248, 276, 314, 328, 358, 376, 416, 430, 469, 486, 531]. Inquiry [477, 106, 122]. Insight [146, 24]. Inspection [32, 483]. Instincts [187]. Institute [285]. Instruments [113]. Insufficient [168]. Integration [511]. Intellectual [490]. Intelligibility [212]. Intensional [523]. Interpretation [332, 193, 198, 11, 221]. Introduction [137, 113, 190, 135]. Introductory [432]. Introspection [32]. Intuition [66]. Intuitive [66]. Inverse [55]. Investigation [311]. Involving [458]. Ions [151]. Irwin [374]. Isaac [216]. Isomorphism [523]. Issues [20].

J [354, 36, 84, 123, 387, 92, 148]. James [51, 372]. Jeffreys [38]. Johannes [81]. John [304, 36, 51, 374, 385, 386]. Johnson [391, 394]. Josef [18].

Justus [22].

Kallen [510]. Kant [18, 333]. Karl [19]. Kinds [61, 4]. Kinship [476]. Klein [138]. Knowledge [79, 419, 217, 394, 258, 402, 222].

L [526, 466, 325, 464, 68, 83]. Language [478, 477, 171, 325, 105]. Laski [292]. Last [71]. Law [266, 487, 74]. Lawrence [482, 17]. Laws [480, 458, 145]. Learning [403]. Leisy [511]. Lenin [454]. Leonard [524]. Letter [322]. Levels [278]. Liberal [510]. Liberalism [471]. Life [300, 235, 393, 51, 390]. Light [258, 364]. Limitations [364]. Limits [156]. Line [71]. Linear [201]. Linguistic [381]. Linsky [524]. Linton [307]. Linus [89]. Liouville [111]. Literati [47]. Living [445, 405, 196, 246]. Locus [412]. Logic [453, 323, 288, 357, 159, 457, 219, 442, 463, 135, 258, 254, 319, 201, 48, 341, 484, 93, 466]. Logical [116, 159, 476, 183, 266]. Logician [317]. Logico [329]. Logico-Historical [329]. Logics [110]. Luther [137]. Lychnos [35]. Lyman [51, 384]. Lynn [124].

M [418, 510, 82, 127, 92]. Mack [306, 371]. Magic [124]. Magnetism [133]. Maier [18]. Making [21]. Malisoff [418]. Man [21, 307, 383, 387]. Mankind [321]. Mann [52]. March [16]. Marcianum [81]. Marshall [105]. Martin [394, 372]. Masquerade [236]. Mass [275]. Mass-Time [275]. Material [76, 128]. Materialism [427, 489, 281]. Mathematical [420, 298, 484, 269, 90, 412, 93]. Mathematics [12, 255, 138, 136]. Matila [390]. Matter [23, 39, 54, 70, 95, 109, 130, 150, 166, 181, 194, 204, 167, 182, 195, 205, 214, 227, 239, 249, 215, 228, 240, 250, 257, 264, 270, 277, 283, 294, 308, 315, 326, 338, 346, 359, 327, 339, 347, 360, 375, 397, 410, 417, 467, 429, 444, 456, 470, 485, 500, 516, 532, 57, 405]. Max [389]. Mayer [92, 92]. Mead [211]. Meaning [100, 78, 106, 122]. Meaningfulness [318]. Meanings [100, 110]. Means [355]. Means-Centering [355]. Measure [413, 301, 423]. Measurement [491, 517]. Mechanical [297]. Mechanics [517, 518, 519, 6, 92]. Mechanism [235, 56, 329, 6]. Mechanisms [404]. Mechanistic [424]. Medical [280]. Melancholie [37]. Membrane [87]. Memorandum [501]. Memoriam [418]. Mensurational [508]. Mental [404]. Metaphysical [477]. Metaphysics [216, 116, 457, 380, 350, 349, 67]. Method [216, 296, 28, 217, 50, 269, 174, 244, 507, 466]. Méthode [66]. Methodological [31]. Methodology [411, 265, 425, 135, 144]. Methods [103, 482]. Michael [528]. Microcoulomb [132]. Middle [74]. Mildred [51]. Miller [464, 86]. Million [255]. Mills [389]. Milton [19]. Mind [146, 423, 218, 104, 320, 16]. Mind-Types [146]. Minds [59]. Models [420, 26, 313]. Modern [120, 457, 55, 510, 366, 302, 450, 67, 8, 12, 282, 364, 48, 21, 383]. Modes [140]. Molecules [151]. Monad [1]. Monism [361]. Moral [19]. Morgenstern [304]. Morgenthau [387]. Morning [492]. Morphological [405]. Motion [12]. Movement [503, 34]. Mr. [460, 49]. Much [396]. Multi [110]. Multi-Valued [110]. Munitz [19]. Murphy [374]. Mysterious [262]. Mystification [75]. Mythology [261].

N [253, 380, 127, 388]. Natural [502, 287, 265, 433, 438, 73]. Natural-Science [265]. Naturalism [329]. Naturalistic [522]. Nature [480, 379, 52, 89, 401, 189, 30, 320, 84, 370]. Nebulae [394]. Necessary [198]. Need [367, 434, 438]. Needs [433]. Neumann [304]. Neurath [21]. Neurodynamics [247]. Neurosis [473]. Neutral [151]. Neutrino [152]. News [443, 455, 468, 515, 529, 284]. Newton [216]. Nicolas [90]. Nilsson [82]. Non [26, 297, 333, 196, 316, 474]. Non-Euclideans [333]. Non-Existence [316]. Non-Formal [474]. Non-Living [196]. Non-Mechanical [297]. Non-Physical [26]. Nonextensional [465]. Norbert [499]. Norman [383]. Normative [286, 292]. Norms [141]. Northrop [453]. Notation [201]. Note [494, 522, 302, 450, 407, 378, 111]. Notes [232, 263, 523]. Notices [443, 455, 468, 515, 529]. Nuance [146]. Null [262, 220, 293]. Number [178, 118].

O [83]. Object [185]. Objectives [103, 364]. Objectivity [64]. Objects [345, 58]. Observed [76]. Obsolete [383]. Obvious [237]. Open [322, 452]. Operationalism [199]. Operationally [42]. Operationism [42, 274]. Operations [343]. Opposition [324]. Organisers [134]. Organism [445, 345]. Organisms [246]. Organization [268, 286, 267, 356]. Origin [300, 60]. Origines [66]. Oskar [304]. Other [59]. Otto [21]. Our [59, 258]. Outline [266]. Outlook [188]. Overlook [237].

P [512, 82, 189]. Pannekoek [454]. Pap [373]. Paper [460]. Paradox [491, 523]. Particulars [9]. past [362]. Paul [51]. Pauling [89]. Peirce [22, 352, 7, 28, 8]. People [336, 391]. Permeability [87]. Personality [413]. Perspective [218]. Pertaining [377]. Pfahler [137]. Pfeiffer [51]. Phenomena [193, 404]. Phenomenology [73]. Philosopher [454]. Philosophical [514, 449, 409, 271, 459, 305]. Philosophies [309, 177]. Philosophy [332, 496, 251, 471, 498, 173, 342, 273, 345, 481, 212, 253, 382, 475, 3, 2, 19, 107, 97, 25, 156, 238, 340, 155, 305, 446, 71, 36, 176, 386, 527, 148]. Philosphies [372]. Phylobiological [247]. Physical [165, 193, 298, 26, 377, 179, 439, 373, 520, 435, 434, 197, 142, 153]. Physicalistic [26]. Physics [11, 210, 56, 350, 421, 340, 145, 15]. Physiological [331]. Picture [298, 51]. Pierre [66, 147]. Pioneer [8]. Place [457, 472]. Planned [385]. Planning [279]. Pledge [530]. Point [426]. Polanyi [528]. Polar [146]. Policy [502, 501, 506]. Politics [387]. Popper [452]. Popular [82]. Popularization [284]. Populations [88]. Positive [151]. Positivism [66, 101, 291]. Positivistic [11]. Possibility [403]. Possible [309]. Postulates [296, 422, 139]. Postulations [317]. Power [387, 392, 512]. Practice [415]. Pragmatic [493, 28]. Pragmatics [409, 451, 495, 494, 399]. Pragmatist [259]. Pratt [48]. Predicate [379]. Prediction [63]. Present [211]. Presentative [202]. Presupposition [64]. Prime [178]. Primitive [46]. Principle [168, 271, 111]. Principles [159, 520]. Priori [373, 72]. Private [5]. Probability [295, 296, 297, 396, 114, 55, 245, 102, 403, 38]. Problem [207, 162, 492, 161, 355, 521, 311, 169, 145]. Problem-Centering [355]. Problems [31, 510, 131, 12]. Process [517]. Processes [421]. Profession [25]. Professor [480, 479]. Program [496]. Progress [411]. Progressive [378]. Projective [482]. Promise [83]. Propositions [324, 117, 33]. Prospects [36]. Protoplasmic [112]. Protozoa [127]. Psychic [246, 267]. Psychical [365, 428, 407]. Psychoanalyse [65]. Psychoanalysis [103]. Psychological [409, 203, 20]. Psychologists [237]. Psychology [31, 57, 504, 103, 15, 174, 331, 48, 526]. Psychophysical [162]. Psychotherapy [342]. Public [5]. Pure [399]. Purpose [424, 235, 3, 208, 172].

Quality [50, 12]. Quandaries [391]. Quantity [12]. Quantum [517, 518, 519, 520]. Quest [321]. Question [361]. Questions [318]. Quine [93].

R [126, 370, 385, 91, 452, 81, 203, 158]. R-Being [158]. R-Systems [203]. Radiation [362]. Ralph [307]. Ramsperger [177]. Rashevsky [90]. Rational [171]. Readings [514]. Realism [348, 199]. Reality [518, 519, 44, 45, 412, 105, 368]. Reason [168]. Received [149, 164, 175, 191, 226]. Reconsidered [480]. Reflections [498]. Reflexes [187]. Reiser [83]. Rejoinder [462]. Relation [98, 45]. Relations [123]. Relativism [419]. Relativity [340, 290]. Relevance [459]. Religion [82, 372, 107]. Remarks [432, 192, 348]. Replacement [424]. Reply [524, 495, 479, 464]. Research [278, 411, 483, 501, 280, 127]. Resolution [521]. Resources [34]. Responsibility [437]. Retardation [53]. Retribution [143]. Review [526, 511, 453, 454, 482, 499, 513, 514, 466, 498, 512, 105, 484, 325, 528, 530, 510, 65, 66, 304, 481, 15, 16, 17, 18, 19, 20, 21, 22, 35, 36, 37, 38, 49, 50, 53, 67, 68, 82, 83, 84, 85, 86, 107, 123, 124, 125, 126, 127, 128, 176, 177, 306, 307, 368, 369, 370, 371, 372, 373, 374, 383, 384, 385, 386, 387, 388, 389, 390, 391, 392, 393, 394, 395, 89, 527, 189, 190, 90, 91, 92, 93, 106, 134, 135, 136, 137, 138, 148, 497, 525, 452, 122, 305, 48]. Reviews [483, 51, 52, 81]. Revolt [47]. Richard [128]. Right [369]. Rigidity [165, 142, 153]. Ritchie [498]. Robert [20, 306, 371]. Rolbiecki [36]. Role [501, 313, 234]. Royal [525]. Rudolf [190]. Runes [176]. Russell [106, 233, 322, 122, 305].

S [453, 20, 8]. Salvaging [331]. Sampling [483]. Santayana [19]. save [462, 461]. Schema [162]. Schroedinger [393]. Science [496, 420, 251, 471, 502, 503, 225, 411, 5, 437, 34, 408, 14, 261, 422, 427, 489, 457, 398, 171, 47, 366, 505, 265, 98, 284, 450, 2, 462, 307, 309, 336, 355, 501, 160, 230, 527, 349, 435, 313, 238, 27, 506, 157, 144, 281, 415, 507, 461, 155, 364, 73, 86, 123, 124, 177, 148, 528, 530, 51, 107, 384, 385]. Sciences [511, 453, 278, 439, 63, 435, 438, 135]. Scientific [513, 140, 466, 192, 188, 217, 83, 256, 367, 401, 425, 343, 279, 285, 521, 174, 487, 244, 311, 41, 387]. Scientist [433]. Scientists [434, 525]. Scope [364]. Scotus [81]. Searles [466]. Sears [51]. Secondary [428]. Secret [428, 316]. Seen [364]. Selection [403]. Selectivity [170]. Self [33]. Self-Contradictory [33]. Sellars [514]. Semantics [480, 260, 190]. Semiotic [312, 459]. Sensation [179]. Sense [381, 258]. Sensum [219]. Sentences [184]. Set [422]. Sextus [120]. Shadow [447]. Sharp [415]. Sherwood [16]. Shewhart [50]. Sidney [374]. Siebold [81]. Sigerist [395]. Significance [41]. Signs [242, 202, 243]. Silence [252]. Similitudes [341]. since [530]. Sir [498, 216]. Skin [71]. Smaller [132]. ’Social [200, 511, 503, 473, 437, 504, 291, 505, 146, 63, 123, 501, 433, 435, 434, 238, 438, 506, 144, 415, 507]. Society [472, 452, 528, 525]. Sociological [366, 145]. Sociologists [402]. Sociology [439, 419, 222, 446, 389]. Solar [497]. Some [449, 31, 323, 11, 115, 441, 523, 12, 433]. Somerville [386]. Soul [85]. Sound [119, 509]. Source [388]. Soviet [489, 281, 386]. Spaces [203]. Specifying [343]. Spending [51]. Spirit [510]. Standpoint [138]. Star [492]. State [385]. Statements [198, 401]. Statistical [483, 493, 50, 91, 92]. Statisticians [55]. Statistics [451, 495, 494]. Status [312]. Steiger [527]. Stevens [15]. Stevenson [325]. Stimson [525]. Storer [460]. Stromberg [85]. Structure [125]. Studies [442, 463]. Study [473, 185, 490, 218, 260, 329, 244, 321, 113]. Subject [57, 273]. Subjectivity [475]. Summers [127]. Survival [274]. Susan [17]. Syllogism [488]. Syllogistic [201]. Symbol [119, 60]. Symbolism [45]. Symposium [511]. Syntactical [180]. Synthesis [161]. System [147, 46, 497]. Systems [203, 159, 196].

T [530, 52, 136]. Table [131]. Tarski [135]. Taxonomic [450]. Taylor [16]. Technical [68]. Techniques [483]. Technology [171]. Teleology [208, 77]. Ten [251]. Tend [237]. Tendencies [101]. Term [200]. Terms [26]. Them [458]. Theorem [111]. Theoretical [441, 224]. Theories [41]. Theory [290, 362, 377, 353, 480, 352, 295, 296, 297, 179, 219, 259, 302, 310, 242, 300, 115, 520, 243, 104, 282, 373, 304, 38]. There [133]. Thermodynamics [131, 189, 91]. These [51]. Thinking [76, 369]. Thorndike [124]. Thorstein [490]. Thought [58, 72]. Thoughts [433]. Three [363]. Thurston [125]. Time [275, 394]. Times [67]. Timothy [37]. Tindall [17]. Tompkins [49]. Trade [503]. Trade-Union [503]. Treatise [37]. Treatment [7]. Triangle [275]. True [133]. Truth [117, 259, 447, 253, 106, 122]. Tweney [68]. Two [478, 4]. Types [146, 341].

Umbrellaology [144]. Uncertainty [111]. Underlying [317]. Understanding [258]. Unification [351]. Unified [193]. Union [503].

Unique [420, 446]. Unity [34]. Universals [9]. Universe [85]. University [395]. Unobservable [44]. Unresting [126]. Urban [105]. Usages [200]. Use [256]. Ushenko [512].

V [93, 138]. Vagueness [14]. Validity [420]. Valuation [265]. Value [502, 398, 414, 265, 223, 104, 319]. Value-Policy [502]. Valued [110]. Values [46, 61]. Varieties [351]. Veblen [490]. Vegetation [185]. Verifiability [117, 61, 29]. Verification [9]. ’Verified [379]. versus [203, 403, 25, 233]. Viewpoint [50]. Virtue [79]. Vital [267]. Vitalism [329]. Vocation [25]. Volume [69, 154, 213, 229, 248, 276, 314, 328, 358, 376, 416, 430, 469, 486, 531]. vs [493, 387, 355, 311].

W [50, 52, 67, 126, 389, 527, 189, 93]. Waddington [513, 134]. Walter [50]. War [173]. Ward [52]. Wave [52]. Way [51, 521]. Weber [389]. Wendell [391]. Werkmeister [480]. West [521]. Wheeler [362]. Which [51]. Whitehead [273, 405, 212, 253, 310, 344, 380, 334]. Who [51]. Wholeness [445]. Wholism [241]. Wieman [388]. Wiener [499]. Wilbur [105]. Wilfrid [514]. William [418, 368]. Winthrop [302]. Wisdom [344, 392]. Wit [344]. without [75, 458]. Wm [17]. Wonderland [49]. Woodbridge [84]. Woodworth [20]. Work [436]. World [307, 186, 197, 29, 510]. Writings [292].

Years [251]. York [17].

Ziporyn [53].

References

Malisoff:1940:WM

[1] William Marias Malisoff. What is a monad? Philosophy of Science, 7(1): 1–6, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

     (electronic). URL http://www.jstor.org/stable/184455.

Lee:1940:PS

[2] Otis Lee. Philosophy and science. Philosophy of Science, 7(1):7–17, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

     tronic). URL http://www.jstor.org/stable/184456.

Lapan:1940:PP

[3] Arthur Lapan. The purpose of philosophy. Philosophy of Science, 7(1):18– 25, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

     (electronic). URL http://www.jstor.org/stable/184457.

Miller:1940:TKC

[4] David L. Miller. Two kinds of certainty. Philosophy of Science, 7(1):26– 35, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184458.

Bridgman:1940:SPP

[5] P. W. Bridgman. Science: Public or private? Philosophy of Science, 7(1):36–48, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184459.

Turner:1940:DBM

[6] J. E. Turner. The distinction between “mechanics” and “mechanism”. Philosophy of Science, 7(1):49–55, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184460.

Goudge:1940:PTI

[7] Thomas A. Goudge. Peirce’s treatment of induction. Philosophy of Science, 7(1):56–68, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184461.

Nagel:1940:CPP

[8] Ernest Nagel. Charles S. Peirce, pioneer of modern empiricism. Philosophy of Science, 7(1):69–80, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184462.

Waters:1940:PUV

[9] Bruce Waters. Particulars, universals and verification. Philosophy of Science, 7(1):81–91, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184463.

Lafleur:1940:FF

[10] Laurence J. Lafleur. The fluxive fallacy. Philosophy of Science, 7(1):92– 96, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184464.

Bushkovitch:1940:SCP

[11] A. V. Bushkovitch. Some consequences of the positivistic interpretation of physics. Philosophy of Science, 7(1):97–102, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184465.

Niebyl:1940:MMS

[12] Karl H. Niebyl. Modern mathematics and some problems of quantity, quality, and motion in economic analysis. Philosophy of Science, 7(1):103– 120, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184466.

Zilsel:1940:HBE

[13] Edgar Zilsel. History and biological evolution. Philosophy of Science, 7 (1):121–128, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184467.

Dexter:1940:SV

[14] Lewis A. Dexter and A. Cornelius Benjamin. Science and vagueness. Philosophy of Science, 7(1):129–131, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184468.

M:1940:BRBa

[15] W. M. M. Book review: The Psychology of Physics by Blamey Stevens. Philosophy of Science, 7(1):132, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184469.

M:1940:BRBb

[16] W. M. M. Book review: The March of Mind by F. Sherwood Taylor. Philosophy of Science, 7(1):132, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184470.

M:1940:BRBc

[17] W. M. M. Book review: D. H. Lawrence and Susan His Cow by Wm. York Tindall. Philosophy of Science, 7(1):132, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184471.

M:1940:BRBd

[18] W. M. M. Book review: On Hegel’s Critique of Kant by Josef Maier. Philosophy of Science, 7(1):133, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184472.

M:1940:BRBe

[19] W. M. M. Book review: The Moral Philosophy of Santayana by Milton Karl Munitz. Philosophy of Science, 7(1):133, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184473.

M:1940:BRBf

[20] W. M. M. Book review: Psychological Issues by Robert S. Woodworth. Philosophy of Science, 7(1):133, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184474.

M:1940:BRBg

[21] W. M. M. Book review: Modern Man in the Making by Otto Neurath. Philosophy of Science, 7(1):133–134, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184475.

M:1940:BRBh

[22] W. M. M. Book review: Charles Peirce’s Empiricism by Justus Buchler. Philosophy of Science, 7(1):134, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184476.

Anonymous:1940:FMa

[23] Anonymous. Front matter. Philosophy of Science, 7(1):??, January 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184454.

Malisoff:1940:WI

[24] William Marias Malisoff. What is insight? Philosophy of Science, 7(2): 135–139, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184803.

Miller:1940:VVP

[25] Cecil H. Miller. Vocation versus profession in philosophy. Philosophy of Science, 7(2):140–150, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184804.

Bergmann:1940:PMN

[26] Gustav Bergmann. On physicalistic models of non-physical terms. Philosophy of Science, 7(2):151–158, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184805.

Sharpe:1940:CAC

[27] Philip B. Sharpe. A critical analysis of the canons of science. Philosophy of Science, 7(2):159–167, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184806.

Hill:1940:PPM

[28] W. H. Hill. Peirce’s “pragmatic” method. Philosophy of Science, 7(2):168– 181, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184807.

Will:1940:VEW

[29] Frederick L. Will. Verifiability and the external world. Philosophy of Science, 7(2):182–191, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184808.

Winn:1940:NC

[30] Ralph B. Winn. The nature of causation. Philosophy of Science, 7(2):192– 204, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184809.

Bergmann:1940:SMPa

[31] Gustav Bergmann. On some methodological problems of psychology. Philosophy of Science, 7(2):205–219, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184810.

Wood:1940:II

[32] Ledger Wood. Inspection and introspection. Philosophy of Science, 7(2): 220–228, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184811.

Lazerowitz:1940:SCP

[33] M. Lazerowitz. Self-contradictory propositions. Philosophy of Science, 7(2):229–240, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184812.

Byrne:1940:EAR

[34] Lee Byrne. An educational application of resources of the unity of science movement. Philosophy of Science, 7(2):241–262, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184813.

M:1940:BRBi

[35] W. M. M. Book review: Lychnos. Philosophy of Science, 7(2):263, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184814.

M:1940:BRBj

[36] W. M. M. Book review: The Prospects of Philosophy by John J. Rolbiecki. Philosophy of Science, 7(2):263, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184815.

M:1940:BRBk

[37] W. M. M. Book review: A Treatise of Melancholie by Timothy Bright. Philosophy of Science, 7(2):263, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184816.

M:1940:BRBl

[38] W. M. M. Book review: Theory of Probability by Harold Jeffreys. Philosophy of Science, 7(2):263–264, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184817.

Anonymous:1940:FMb

[39] Anonymous. Front matter. Philosophy of Science, 7(2):??, April 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184802.

Malisoff:1940:WF

[40] William Marias Malisoff. What is freedom? Philosophy of Science, 7(3): 265–272, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184845.

Swann:1940:SST

[41] W. F. G. Swann. The significance of scientific theories. Philosophy of Science, 7(3):273–287, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184846.

Hart:1940:OAO

[42] Hornell Hart. Operationism analysed operationally. Philosophy of Science, 7(3):288–313, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184847.

Lillie:1940:BC

[43] Ralph S. Lillie. Biological causation. Philosophy of Science, 7(3):314– 336, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184848.

Metcalf:1940:RU

[44] W. V. Metcalf. The reality of the unobservable. Philosophy of Science, 7 (3):337–341, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184849.

Sisson:1940:RRS

[45] Edward O. Sisson. Relation in reality and symbolism. Philosophy of Science, 7(3):342–354, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184850.

Lee:1940:PSV

[46] D. Demetracopoulou Lee. A primitive system of values. Philosophy of Science, 7(3):355–378, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184851.

Goran:1940:LRA

[47] Morris Goran. The literati revolt against science. Philosophy of Science, 7(3):379–384, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184852.

Wood:1940:BRB

[48] Ledger Wood. Book review: The Logic of Modern Psychology by Carroll C. Pratt. Philosophy of Science, 7(3):385–386, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184853.

M:1940:BRBm

[49] W. M. M. Book review: Mr. Tompkins in Wonderland by G. Gamow. Philosophy of Science, 7(3):386, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184854.

M:1940:BRBn

[50] W. M. M. Book review: Statistical Method from the Viewpoint of Quality Control by Walter A. Shewhart; W. E. Deming. Philosophy of Science, 7 (3):386, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184855.

M:1940:BRBo

[51] W. M. M. Book reviews: Science in Your Life by John Pfeiffer; Picture of Health by James Clarke; Getting and Spending by Mildred Adams; Who Are These Americans? by Paul B. Sears; Which Way America? by Lyman Bryson. Philosophy of Science, 7(3):386–387, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184856.

M:1940:BRBp

[52] W. M. M. Book reviews: The Nature of the Atom by G. K. T. Conn; The Nature of Crystals by A. G. Ward; The Wave Nature of the Electron by G. K. T. Conn; The Cyclotron by W. B. Mann. Philosophy of Science, 7 (3):387–388, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184857.

M:1940:BRBq

[53] W. M. M. Book review: Cosmo-Retardation by I. Ziporyn. Philosophy of Science, 7(3):388, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184858.

Anonymous:1940:FMc

[54] Anonymous. Front matter. Philosophy of Science, 7(3):??, July 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184844.

Gordon:1940:IPM

[55] Robert Dean Gordon. Inverse probability and modern statisticians. Philosophy of Science, 7(4):389–399, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184538.

Malisoff:1940:PDM

[56] William Marias Malisoff. Physics: The decline of mechanism. Philosophy of Science, 7(4):400–414, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184539.

Bergmann:1940:SMPb

[57] Gustav Bergmann. The subject matter of psychology. Philosophy of Science, 7(4):415–433, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184540.

Melden:1940:TO

[58] A. I. Melden. Thought and its objects. Philosophy of Science, 7(4):434– 441, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184541.

Dotterer:1940:OCO

[59] Ray H. Dotterer. Our certainty of other minds. Philosophy of Science, 7 (4):442–450, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184542.

White:1940:SOB

[60] Leslie A. White. The symbol: The origin and basis of human behavior. Philosophy of Science, 7(4):451–463, October 1940. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184543.

Lepley:1940:VDK

[61] Ray Lepley. The verifiability of different kinds of facts and values. Philosophy of Science, 7(4):464–475, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184544.

Ritchie:1940:FEE

[62] Benbow F. Ritchie and Abraham Kaplan. A framework for an empirical ethics. Philosophy of Science, 7(4):476–491, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184545.

Kaplan:1940:PSS

[63] Oscar Kaplan. Prediction in the social sciences. Philosophy of Science, 7 (4):492–498, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184546.

Ratner:1940:POH

[64] Sidney Ratner. Presupposition and objectivity in history. Philosophy of Science, 7(4):499–505, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184547.

Harms:1940:BRBa

[65] Ernst Harms. Book review: La Psychoanalyse by Charles Baudouin. Philosophy of Science, 7(4):506, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184548.

Harms:1940:BRBb

[66] Ernst Harms. Book review: Essai sur les Origines Intuitive du Positivism; Méthode et Intuition Chez Auguste Comte by Pierre Ducasse. Philosophy of Science, 7(4):506, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184549.

M:1940:BRBr

[67] W. M. M. Book review: Metaphysics in Modern Times by D. W. Gotshalk. Philosophy of Science, 7(4):507, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184550.

M:1940:BRBs

[68] W. M. M. Book review: Chambers’s Technical Dictionary by C. F. Tweney; L. E. C. Hughes. Philosophy of Science, 7(4):507, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184551.

Anonymous:1940:VI

[69] Anonymous. Volume information. Philosophy of Science, 7(4):509–511, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184536.

Anonymous:1940:FMd

[70] Anonymous. Front matter. Philosophy of Science, 7(4):??, October 1940. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184537.

Bentley:1941:HSP

[71] Arthur F. Bentley. The human skin: Philosophy’s last line of defense. Philosophy of Science, 8(1):1–19, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184360.

Miller:1941:PCT

[72] David L. Miller. The a priori in contemporary thought. Philosophy of Science, 8(1):20–25, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184361.

Zilsel:1941:PNS

[73] Edgar Zilsel. Phenomenology and natural science. Philosophy of Science, 8(1):26–32, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184362.

Toms:1941:LEM

[74] Eric Toms. The law of excluded middle. Philosophy of Science, 8(1):33– 38, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184363.

Malisoff:1941:CEM

[75] William Marias Malisoff. Chemistry: Emergence without mystification. Philosophy of Science, 8(1):39–52, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184364.

Bateson:1941:ETA

[76] Gregory Bateson. Experiments in thinking about observed ethnological material. Philosophy of Science, 8(1):53–68, January 1941. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184365.

Weber:1941:TE

[77] Alden O. Weber and David Rapaport. Teleology and the emotions. Philosophy of Science, 8(1):69–82, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184366.

Krikorian:1941:MB

[78] Y. H. Krikorian. Meaning as behavior. Philosophy of Science, 8(1):83–88, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184367.

Fries:1941:VK

[79] Horace S. Fries. Virtue is knowledge. Philosophy of Science, 8(1):89–99, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184368.

Brightman:1941:HBE

[80] Edgar Sheffield Brightman and Edgar Zilsel. History and biological evolution. Philosophy of Science, 8(1):100–101, January 1941. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184369.

W:1941:BRB

[81] W. Book reviews: The Astronomy of Johannes Scotus Erigena by E. von Erhardt-Siebold; R. von Erhardt; Cosmology in the “Annotationes in Marcianum” by E. von Erhardt-Siebold; R. von Erhardt. Philosophy of Science, 8(1):102, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184370.

M:1941:BRBa

[82] M. Book review: Greek Popular Religion by M. P. Nilsson. Philosophy of Science, 8(1):102–103, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184371.

M:1941:BRBb

[83] W. M. M. Book review: The Promise of Scientific Humanism by O. L. Reiser. Philosophy of Science, 8(1):103, January 1941. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184372.

M:1941:BRBc

[84] W. M. M. Book review: An Essay on Nature by F. J. E. Woodbridge. Philosophy of Science, 8(1):103–104, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184373.

M:1941:BRBd

[85] W. M. M. Book review: The Soul of the Universe by Gustaf Stromberg. Philosophy of Science, 8(1):104, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184374.

M:1941:BRBe

[86] M. Book review: History and Science by Hugh Miller. Philosophy of Science, 8(1):104, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184375.

Reiner:1941:DBM

[87] John M. Reiner. Diffusion and biological membrane permeability. II. Philosophy of Science, 8(1):105–114, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184376.

Avrami:1941:GDP

[88] Melvin Avrami. Geometry and dynamics of populations. Philosophy of Science, 8(1):115–132, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184377.

Malisoff:1941:BRB

[89] William Marias Malisoff. Book review: The Nature of the Chemical Bond by Linus Pauling. Philosophy of Science, 8(1):133, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184378.

Reiner:1941:BRBa

[90] John M. Reiner. Book review: Advances and Applications of Mathematical Biology by Nicolas Rashevsky. Philosophy of Science, 8(1):133–134, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184379.

Reiner:1941:BRBb

[91] John M. Reiner. Book review: Statistical Thermodynamics by R. H. Fowler; E. A. Guggenheim. Philosophy of Science, 8(1):134–135, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184380.

Reiner:1941:BRBc

[92] John M. Reiner. Book review: Statistical Mechanics by J. E. Mayer; M. G. Mayer. Philosophy of Science, 8(1):135–136, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184381.

Reiner:1941:BRBd

[93] John M. Reiner. Book review: Mathematical Logic by W. V. Quine. Philosophy of Science, 8(1):136, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184382.

Anonymous:1941:Ea

[94] Anonymous. Editorial. Philosophy of Science, 8(1):137, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184383.

Anonymous:1941:FMa

[95] Anonymous. Front matter. Philosophy of Science, 8(1):??, January 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184359.

Anonymous:1941:Eb

[96] Anonymous. Editorial. Philosophy of Science, 8(2):139, April 1941. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184608.

Malisoff:1941:HP

[97] William Marias Malisoff. On having a philosophy. Philosophy of Science, 8(2):140–141, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184609.

Jones:1941:IRS

[98] Philip Chapin Jones. Idealism and its relation to science. Philosophy of Science, 8(2):142–146, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184610.

Bentley:1941:DD

[99] Arthur F. Bentley. Decrassifying Dewey. Philosophy of Science, 8(2):147– 156, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184611.

Gomperz:1941:MM

[100] H. Gomperz. The meanings of “meaning”. Philosophy of Science, 8(2): 157–183, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184612.

Hartshorne:1941:ATP

[101] Charles Hartshorne. Anthropomorphic tendencies in positivism. Philosophy of Science, 8(2):184–203, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184613.

Kemble:1941:PC

[102] Edwin C. Kemble. The probability concept. Philosophy of Science, 8(2): 204–232, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184614.

Kardiner:1941:PPC

[103] A. Kardiner. Psychoanalysis and psychology. A comparison of methods and objectives. Philosophy of Science, 8(2):233–254, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184615.

Norton:1941:TVT

[104] William J. Norton, Jr. Towards a value theory of mind. Philosophy of Science, 8(2):255–263, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184616.

Bross:1941:BRB

[105] John R. Bross and George J. Bowdery. Book review: Language and Reality by Wilbur Marshall Urban. Philosophy of Science, 8(2):264–265, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184617.

Reiner:1941:BRBe

[106] John M. Reiner. Book review: An Inquiry into Meaning and Truth by Bertrand Russell. Philosophy of Science, 8(2):266–267, April 1941. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184618.

M:1941:BRBf

[107] W. M. M. Book review: Science, Philosophy and Religion. Philosophy of Science, 8(2):267, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184619.

Anonymous:1941:Aa

[108] Anonymous. Announcements. Philosophy of Science, 8(2):268, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184620.

Anonymous:1941:FMb

[109] Anonymous. Front matter. Philosophy of Science, 8(2):269, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184607.

Malisoff:1941:MMV

[110] William Marias Malisoff. Meanings in multi-valued logics. Philosophy of Science, 8(2):271–274, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184621.

VanVleck:1941:NLT

[111] J. H. Van Vleck. Note on Liouville’s Theorem and the Heisenberg Uncertainty Principle. Philosophy of Science, 8(2):275–279, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184622.

Heilbrunn:1941:PA

[112] L. V. Heilbrunn. Protoplasmic activity. Philosophy of Science, 8(2):280– 286, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184623.

Weiss:1941:ISI

[113] Paul Weiss. An introduction to a study of instruments. Philosophy of Science, 8(3):287–296, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184301.

Dotterer:1941:IEP

[114] Ray H. Dotterer. Ignorance and equal probability. Philosophy of Science, 8(3):297–303, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184302.

Keeton:1941:SAT

[115] Morris T. Keeton. Some ambiguities in the theory of the conservation of energy. Philosophy of Science, 8(3):304–319, July 1941. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184303.

Erickson:1941:MLE

[116] Ralph W. Erickson. The metaphysics of a logical empiricist. Philosophy of Science, 8(3):320–328, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184304.

Ducasse:1941:TVP

[117] C. J. Ducasse. Truth, verifiability, and propositions about the future. Philosophy of Science, 8(3):329–337, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184305.

Ushenko:1941:CN

[118] A. P. Ushenko. Class and number. Philosophy of Science, 8(3):338–351, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184306.

Nock:1941:SS

[119] S. A. Nock. Sound and symbol. Philosophy of Science, 8(3):352–370, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184307.

Chisholm:1941:SEM

[120] Roderick M. Chisholm. Sextus empiricus and modern empiricism. Philosophy of Science, 8(3):371–384, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184308.

Harding:1941:EC

[121] T. Swann Harding. Exploitation of the creators. Philosophy of Science, 8(3):385–390, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184309.

Ushenko:1941:BRB

[122] A. P. Ushenko. Book review: An Inquiry into Meaning and Truth by B. Russell. Philosophy of Science, 8(3):391–392, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184310.

M:1941:BRBg

[123] W. M. M. Book review: The Social Relations of Science by J. G. Crowther. Philosophy of Science, 8(3):392, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184311.

M:1941:BRBh

[124] W. M. M. Book review: History of Magic and Experimental Science by Lynn Thorndike. Philosophy of Science, 8(3):393, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184312.

M:1941:BRBi

[125] W. M. M. Book review: The Structure of Art by Carl Thurston. Philosophy of Science, 8(3):393, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184313.

M:1941:BRBj

[126] W. M. M. Book review: Unresting Cells by R. W. Gerard. Philosophy of Science, 8(3):393–394, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184314.

M:1941:BRBk

[127] W. M. M. Book review: Protozoa in Biological Research by G. N. Calkins; F. M. Summers. Philosophy of Science, 8(3):394, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184315.

M:1941:BRBl

[128] W. M. M. Book review: The Material Basis of Evolution by Richard Goldschmidt. Philosophy of Science, 8(3):394–395, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184316.

Anonymous:1941:Ab

[129] Anonymous. Announcements. Philosophy of Science, 8(3):396, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184317.

Anonymous:1941:FMc

[130] Anonymous. Front matter. Philosophy of Science, 8(3):397, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184300.

Malisoff:1941:PTE

[131] William Marias Malisoff. The problems of thermodynamics. I. The energy table. Philosophy of Science, 8(3):399–402, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184318.

Ehrenhaft:1941:MEC

[132] Felix Ehrenhaft. The microcoulomb experiment. Charges smaller than the electronic charge. Philosophy of Science, 8(3):403–457, July 1941.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184319.

Ehrenhaft:1941:TTM

[133] Felix Ehrenhaft and Leo Banet. Is there “true magnetism” or not? Philosophy of Science, 8(3):458–462, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184320.

Reiner:1941:BRBf

[134] John M. Reiner. Book review: Organisers and Genes by C. H. Waddington. Philosophy of Science, 8(3):463, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184321.

Reiner:1941:BRBg

[135] John M. Reiner. Book review: Introduction to Logic and to the Methodology of Deductive Sciences by Alfred Tarski. Philosophy of Science, 8(3): 463–464, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184322.

Reiner:1941:BRBh

[136] John M. Reiner. Book review: The Development of Mathematics by E. T. Bell. Philosophy of Science, 8(3):464–465, July 1941. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184323.

Reiner:1941:BRBi

[137] John M. Reiner. Book review: An Introduction to Differential Geometry by Luther Pfahler Eisenhart. Philosophy of Science, 8(3):465, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184324.

Reiner:1941:BRBj

[138] John M. Reiner. Book review: Elementary Mathematics from an Advanced Standpoint, v. I by Felix Klein. Philosophy of Science, 8(3):466, July 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184325.

Malisoff:1941:PE

[139] William Marias Malisoff. On the postulates of empiricism. Philosophy of Science, 8(4):467–485, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184509.

Benjamin:1941:MSE

[140] A. Cornelius Benjamin. Modes of scientific explanation. Philosophy of Science, 8(4):486–492, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184510.

Bowdery:1941:CN

[141] George J. Bowdery. Conventions and norms. Philosophy of Science, 8 (4):493–505, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184511.

Weinberg:1941:RFPa

[142] Carlton B. Weinberg. Rigidity, force and physical geometry. Philosophy of Science, 8(4):506–532, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184512.

Kelsen:1941:CR

[143] Hans Kelsen. Causality and retribution. Philosophy of Science, 8(4):533– 556, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184513.

Somerville:1941:UMS

[144] John Somerville. Umbrellaology, or, methodology in social science. Philosophy of Science, 8(4):557–566, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184514.

Zilsel:1941:PPH

[145] Edgar Zilsel. Physics and the problem of historico-sociological laws. Philosophy of Science, 8(4):567–579, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184515.

Humphrey:1941:SIN

[146] Norman D. Humphrey. Social insight, nuance, and mind-types: A polar hypothesis. Philosophy of Science, 8(4):580–584, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184516.

Carnus:1941:CSP

[147] Juliette Carnus. The cosmological system of Pierre Bayle. Philosophy of Science, 8(4):585–597, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184517.

Reiser:1941:BRB

[148] Oliver L. Reiser. Book review: Philosophy as a Science by C. J. Ducasse. Philosophy of Science, 8(4):598, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184518.

Anonymous:1941:BR

[149] Anonymous. Books received. Philosophy of Science, 8(4):598–599, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184519.

Anonymous:1941:FMd

[150] Anonymous. Front matter. Philosophy of Science, 8(4):601, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184508.

Margenau:1941:FBP

[151] Henry Margenau. On the forces between positive ions and neutral molecules. Philosophy of Science, 8(4):603–613, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184520.

Stern:1941:NC

[152] Alexander W. Stern. The neutrino concept. Philosophy of Science, 8 (4):614–617, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184521.

Weinberg:1941:RFPb

[153] Carleton B. Weinberg. Rigidity, force and physical geometry. Philosophy of Science, 8(4):618–623, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184522.

Anonymous:1941:VI

[154] Anonymous. Volume information. Philosophy of Science, 8(4):624–627, October 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184507.

Winn:1942:PS

[155] Ralph B. Winn. Philosophy and science. Philosophy of Science, 9(1):1– 18, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184678.

Miller:1942:LFP

[156] Cecil H. Miller. The limits of freedom in philosophy. Philosophy of Science, 9(1):19–29, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184679.

Smith:1942:GCC

[157] Mapheus Smith. The greatest common component in science. Philosophy of Science, 9(1):30–36, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184680.

Lafleur:1942:RB

[158] Laurence J. Lafleur. The R-being. Philosophy of Science, 9(1):37–39, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184681.

Farber:1942:LSP

[159] Marvin Farber. Logical systems and the principles of logic. Philosophy of Science, 9(1):40–54, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184682.

Metcalf:1942:IS

[160] W. V. Metcalf. Idealism and science. Philosophy of Science, 9(1):55–58, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184683.

Lillie:1942:PSB

[161] Ralph S. Lillie. The problem of synthesis in biology. Philosophy of Science, 9(1):59–71, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184684.

Bergmann:1942:ESP

[162] Gustav Bergmann. An empiricist schema of the psychophysical problem. Philosophy of Science, 9(1):72–91, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184685.

Gerard:1942:BBE

[163] R. W. Gerard. A biological basis for ethics. Philosophy of Science, 9 (1):92–120, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184686.

Anonymous:1942:BRa

[164] Anonymous. Books received. Philosophy of Science, 9(1):121, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184687.

Anonymous:1942:ERF

[165] Anonymous. Erratum: Rigidity, force and physical geometry. Philosophy of Science, 9(1):121, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184688.

Anonymous:1942:BMa

[166] Anonymous. Back matter. Philosophy of Science, 9(1):??, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184689.

Anonymous:1942:FMa

[167] Anonymous. Front matter. Philosophy of Science, 9(1):??, January 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184677.

Dubs:1942:PIR

[168] Homer H. Dubs. The principle of insufficient reason. Philosophy of Science, 9(2):123–131, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184421.

Ushenko:1942:PCI

[169] A. P. Ushenko. The problem of causal inference. Philosophy of Science, 9(2):132–138, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184422.

Phillips:1942:CS

[170] Herbert J. Phillips. Causation and selectivity. Philosophy of Science, 9(2):139–145, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184423.

Gerr:1942:LSR

[171] Stanley Gerr. Language and science the rational, functional language of science and technology. Philosophy of Science, 9(2):146–161, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184424.

Weiss:1942:PP

[172] Paul Weiss. The purpose of purpose. Philosophy of Science, 9(2):162– 165, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184425.

Cerf:1942:PW

[173] Walter Cerf. Philosophy and this War. Philosophy of Science, 9(2):166– 182, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184426.

Rogosin:1942:SMC

[174] H. Rogosin. Scientific method in current psychology. Philosophy of Science, 9(2):183–188, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184427.

Anonymous:1942:BRb

[175] Anonymous. Books received. Philosophy of Science, 9(2):189, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184430.

M:1942:BRBa

[176] W. M. M. Book review: The Dictionary of Philosophy by Dagobert D. Runes. Philosophy of Science, 9(2):189, April 1942. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184428.

M:1942:BRBb

[177] W. M. M. Book review: Philosophies of Science by Albert G. Ramsperger. Philosophy of Science, 9(2):189, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184429.

Hartman:1942:PNC

[178] Robert S. Hartman. Prime number and cosmical number. Philosophy of Science, 9(2):190–196, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184431.

Culbertson:1942:PTS

[179] James T. Culbertson. A physical theory of sensation. Philosophy of Science, 9(2):197–226, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184432.

Bergmann:1942:SAC

[180] Gustav Bergmann. Syntactical analysis of the class calculus. Philosophy of Science, 9(2):227–232, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184433.

Anonymous:1942:BMb

[181] Anonymous. Back matter. Philosophy of Science, 9(2):??, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184434.

Anonymous:1942:FMb

[182] Anonymous. Front matter. Philosophy of Science, 9(2):??, April 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184420.

Hester:1942:LDD

[183] William Hester. The logical destiny of data. Philosophy of Science, 9(3): 233–238, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184766.

Waters:1942:BSI

[184] Bruce Waters. Basic sentences and incorrigibility. Philosophy of Science, 9(3):239–244, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184767.

Egler:1942:VOS

[185] Frank E. Egler. Vegetation as an object of study. Philosophy of Science, 9(3):245–260, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184768.

Nelson:1942:EWI

[186] Everett J. Nelson. The external world and induction. Philosophy of Science, 9(3):261–267, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184769.

Weiss:1942:HIR

[187] Paul Weiss. Habits, instincts and reflexes. Philosophy of Science, 9(3): 268–274, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184770.

Feibleman:1942:SOC

[188] James Feibleman. The scientific outlook of Cézanne. Philosophy of Science, 9(3):275–280, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184771.

R:1942:BRBa

[189] J. M. R. Book review: The Nature of Thermodynamics by P. W. Bridgman. Philosophy of Science, 9(3):281, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184772.

R:1942:BRBb

[190] J. M. R. Book review: Introduction to Semantics by Rudolf Carnap. Philosophy of Science, 9(3):281–282, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184773.

Anonymous:1942:BRc

[191] Anonymous. Books received. Philosophy of Science, 9(3):282, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184774.

Bergmann:1942:RCE

[192] Gustav Bergmann. Remarks concerning the epistemology of scientific empiricism. Philosophy of Science, 9(3):283–293, July 1942. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184775.

Baldin:1942:EBC

[193] Siluan F. Baldin. Energy as the basic concept for a unified interpretation of physical phenomena. Philosophy of Science, 9(3):294–305, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184776.

Anonymous:1942:BMc

[194] Anonymous. Back matter. Philosophy of Science, 9(3):??, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184777.

Anonymous:1942:FMc

[195] Anonymous. Front matter. Philosophy of Science, 9(3):??, July 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184765.

Lillie:1942:LSN

[196] Ralph S. Lillie. Living systems and non-living systems. Philosophy of Science, 9(4):307–322, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184691.

Stromberg:1942:CPW

[197] Gustaf Stromberg. Coherence in the physical world. Philosophy of Science, 9(4):323–334, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184692.

Black:1942:CGI

[198] Max Black. Conventionalism in geometry and the interpretation of necessary statements. Philosophy of Science, 9(4):335–349, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184693.

Hartung:1942:OIR

[199] Frank E. Hartung. Operationalism: Idealism or realism? Philosophy of Science, 9(4):350–355, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184694.

Bowdery:1942:UTS

[200] Barbara K. Bowdery. Usages of the term ’social’. Philosophy of Science, 9(4):356–361, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184695.

Thompson:1942:SLL

[201] Samuel M. Thompson. Syllogistic logic in linear notation. Philosophy of Science, 9(4):362–366, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184696.

Moore:1942:IPF

[202] Willis Moore, Gustave Bergmann, and Ray H. Dotterer. The indexical and the presentative functions of signs. Philosophy of Science, 9(4):367– 375, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184697.

Creegan:1942:RSV

[203] Robert Francis Creegan. R-systems versus psychological spaces. Philosophy of Science, 9(4):376–379, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184698.

Anonymous:1942:BMd

[204] Anonymous. Back matter. Philosophy of Science, 9(4):??, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184699.

Anonymous:1942:FMd

[205] Anonymous. Front matter. Philosophy of Science, 9(4):??, October 1942. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184690.

Boodin:1943:CA

[206] John Elof Boodin. Cosmic attributes. Philosophy of Science, 10(1):1–12, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184876.

Benjamin:1943:EPE

[207] A. Cornelius Benjamin. The essential problem of empiricism. Philosophy of Science, 10(1):13–17, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184877.

Rosenblueth:1943:BPT

[208] Arturo Rosenblueth, Norbert Wiener, and Julian Bigelow. Behavior, purpose and teleology. Philosophy of Science, 10(1):18–24, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184878.

Seifriz:1943:CII

[209] William Seifriz. Creative imagination and indeterminism. Philosophy of Science, 10(1):25–33, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184879.

Jones:1943:PI

[210] Philip Chapin Jones. Physics and idealism. Philosophy of Science, 10 (1):34–39, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184880.

Miller:1943:GHM

[211] David L. Miller. G. H. Mead’s conception of “present”. Philosophy of Science, 10(1):40–46, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184881.

Johnson:1943:IWP

[212] A. H. Johnson. The intelligibility of Whitehead’s philosophy. Philosophy of Science, 10(1):47–55, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184882.

Anonymous:1943:VIa

[213] Anonymous. Volume information. Philosophy of Science, 10(1):56–57, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184874.

Anonymous:1943:BMa

[214] Anonymous. Back matter. Philosophy of Science, 10(1):??, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184883.

Anonymous:1943:FMa

[215] Anonymous. Front matter. Philosophy of Science, 10(1):??, January 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184875.

Burtt:1943:MMS

[216] E. A. Burtt. Method and metaphysics in Sir Isaac Newton. Philosophy of Science, 10(2):57–66, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184288.

Lee:1943:SMK

[217] Harold N. Lee. Scientific method and knowledge. Philosophy of Science, 10(2):67–74, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184289.

Miller:1943:MSP

[218] Cecil H. Miller. Mind — a study in perspective. Philosophy of Science, 10 (2):75–80, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184290.

Gentry:1943:LST

[219] George Gentry. The logic of the sensum theory. Philosophy of Science, 10 (2):81–89, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184291.

Carmichael:1943:ANC

[220] Peter A. Carmichael. Animadversion on the null class. Philosophy of Science, 10(2):90–94, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184292.

Gauss:1943:II

[221] Charles E. Gauss. The interpretation of implication. Philosophy of Science, 10(2):95–103, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184293.

Wolff:1943:SKE

[222] Kurt H. Wolff. The sociology of knowledge: Emphasis on an empirical attitude. Philosophy of Science, 10(2):104–123, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184294.

Lepley:1943:IFV

[223] Ray Lepley. The identity of fact and value. Philosophy of Science, 10(2): 124–130, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184295.

Winthrop:1943:TBB

[224] Henry Winthrop. The theoretical basis of biotypology. Philosophy of Science, 10(2):131–139, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184296.

Bateson:1943:SD

[225] Gregory Bateson. The science of decency. Philosophy of Science, 10(2): 140–142, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184297.

Anonymous:1943:BR

[226] Anonymous. Books received. Philosophy of Science, 10(2):143, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184298.

Anonymous:1943:BMb

[227] Anonymous. Back matter. Philosophy of Science, 10(2):??, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184299.

Anonymous:1943:FMb

[228] Anonymous. Front matter. Philosophy of Science, 10(2):??, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184287.

Anonymous:1943:VIb

[229] Anonymous. Volume information. Philosophy of Science, 10(2):??, April 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184286.

Miller:1943:BDS

[230] David L. Miller. Basic decisions in science. Philosophy of Science, 10(3): 145–148, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184900.

Mueller:1943:BB

[231] Gustav E. Mueller. On being and becoming. Philosophy of Science, 10 (3):149–162, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184901.

Bergmann:1943:NI

[232] Gustav Bergmann. Notes on identity. Philosophy of Science, 10(3):163– 166, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184902.

Smart:1943:CVR

[233] Harold R. Smart. Cassirer versus Russell. Philosophy of Science, 10(3): 167–175, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184903.

Witzemann:1943:RCB

[234] Edgar J. Witzemann. The role of catalysis in biological causation. Philosophy of Science, 10(3):176–183, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184904.

Krikorian:1943:LMP

[235] Y. H. Krikorian. Life, mechanism and purpose. Philosophy of Science, 10(3):184–190, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184905.

Nabours:1943:ME

[236] Robert K. Nabours. The masquerade of ESP. Philosophy of Science, 10(3):191–203, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184906.

Ichheiser:1943:WPT

[237] Gustav Ichheiser. Why psychologists tend to overlook certain “obvious” facts. Philosophy of Science, 10(3):204–207, July 1943. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184907.

Schrickel:1943:PSS

[238] H. G. Schrickel. Philosophy of science and social philosophy. Philosophy of Science, 10(3):208–212, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184908.

Anonymous:1943:BMc

[239] Anonymous. Back matter. Philosophy of Science, 10(3):??, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184909.

Anonymous:1943:FMc

[240] Anonymous. Front matter. Philosophy of Science, 10(3):??, July 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184899.

Boodin:1943:AW

[241] John Elof Boodin. Analysis and wholism. Philosophy of Science, 10(4): 213–229, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184267.

Kaplan:1943:CAT

[242] Abraham Kaplan. Content analysis and the theory of signs. Philosophy of Science, 10(4):230–247, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184268.

Morris:1943:CAT

[243] Charles Morris. Content analysis and the theory of signs: Comments. Philosophy of Science, 10(4):247–249, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184269.

Smith:1943:MES

[244] Mapheus Smith. The method of elimination in scientific study. Philosophy of Science, 10(4):250–254, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184270.

Hawkins:1943:EEP

[245] David Hawkins. Existential and epistemic probability. Philosophy of Science, 10(4):255–261, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184271.

Lillie:1943:PFL

[246] Ralph S. Lillie. The psychic factor in living organisms. Philosophy of Science, 10(4):262–270, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184272.

Burrow:1943:NBP

[247] Trigant Burrow. The neurodynamics of behavior. A phylobiological foreword. Philosophy of Science, 10(4):271–288, October 1943. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184273.

Anonymous:1943:VIc

[248] Anonymous. Volume information. Philosophy of Science, 10(4):289–290, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184265.

Anonymous:1943:BMd

[249] Anonymous. Back matter. Philosophy of Science, 10(4):??, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184274.

Anonymous:1943:FMd

[250] Anonymous. Front matter. Philosophy of Science, 10(4):??, October 1943. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184266.

Anonymous:1944:EPS

[251] Anonymous. Editorial: Philosophy of science after ten years. Philosophy of Science, 11(1):1–2, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184885.

Malisoff:1944:CES

[252] William Marias Malisoff. Cratylus or an essay on silence (not illustrated). Philosophy of Science, 11(1):3–8, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184886.

Johnson:1944:TBG

[253] A. H. Johnson. “Truth, beauty and goodness” in the philosophy of A. N. Whitehead. Philosophy of Science, 11(1):9–29, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184887.

Ross:1944:IL

[254] Alf Ross. Imperatives and logic. Philosophy of Science, 11(1):30–46, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184888.

Pavlov:1944:MDM

[255] Vassily Pavlov. Mathematics for the doctor in the million. Philosophy of Science, 11(1):47–52, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184889.

Meadows:1944:SUH

[256] Paul Meadows. The scientific use of historical data. Philosophy of Science, 11(1):53–58, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184890.

Anonymous:1944:FMa

[257] Anonymous. Front matter. Philosophy of Science, 11(1):??, January 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184884.

Ritter:1944:LOC

[258] William E. Ritter. Logic in our common knowledge or logic in the light of common sense, common knowledge, and common understanding. Philosophy of Science, 11(2):59–81, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184353.

Hinshaw:1944:PTT

[259] Virgil G. Hinshaw, Jr. The pragmatist theory of truth. Philosophy of Science, 11(2):82–92, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184354.

Sapir:1944:GSS

[260] Edward Sapir. Grading, a study in semantics. Philosophy of Science, 11(2):93–116, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184355.

Feibleman:1944:MS

[261] James Feibleman. The mythology of science. Philosophy of Science, 11(2): 117–121, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184356.

Black:1944:MNC

[262] Max Black. A mysterious null class. Philosophy of Science, 11(2):122, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184357.

Bergmann:1944:NI

[263] Gustav Bergmann. Notes on identity. Philosophy of Science, 11(2):123– 124, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184358.

Anonymous:1944:FMb

[264] Anonymous. Front matter. Philosophy of Science, 11(2):??, April 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184352.

Hull:1944:VVN

[265] Clark L. Hull. Value, valuation, and natural-science methodology. Philosophy of Science, 11(3):125–141, July 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184602.

Oppenheim:1944:OLA

[266] Felix E. Oppenheim. Outline of a logical analysis of law. Philosophy of Science, 11(3):142–160, July 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184603.

Lillie:1944:VOP

[267] Ralph S. Lillie. Vital organization and the psychic factor. Philosophy of Science, 11(3):161–170, July 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184604.

Dunlop:1944:O

[268] W. R. Dunlop. Organization. Philosophy of Science, 11(3):171–177, July 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184605.

Montague:1944:MID

[269] Harriet F. Montague. The method of infinite descent and the method of mathematical induction. Philosophy of Science, 11(3):178–185, July 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184606.

Anonymous:1944:FMc

[270] Anonymous. Front matter. Philosophy of Science, 11(3):??, July 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184601.

Margenau:1944:EPP

[271] Henry Margenau. The exclusion principle and its philosophical importance. Philosophy of Science, 11(4):187–208, October 1944. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184797.

Bergmann:1944:HHE

[272] Gustav Bergmann. Holism, historicism, and emergence. Philosophy of Science, 11(4):209–221, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184798.

Gentry:1944:SWP

[273] George Gentry. The subject in Whitehead’s philosophy. Philosophy of Science, 11(4):222–226, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184799.

Hartung:1944:OCS

[274] Frank E. Hartung. Operationism as a cultural survival. Philosophy of Science, 11(4):227–232, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184800.

Dodd:1944:MTT

[275] Stuart Carter Dodd. A mass-time triangle. Philosophy of Science, 11 (4):233–244, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184801.

Anonymous:1944:VI

[276] Anonymous. Volume information. Philosophy of Science, 11(4):245, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184795.

Anonymous:1944:FMd

[277] Anonymous. Front matter. Philosophy of Science, 11(4):??, October 1944. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184796.

Bailey:1945:LRB

[278] Orville T. Bailey. Levels of research in the biological sciences. Philosophy of Science, 12(1):1–7, January 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184715.

Rautenstrauch:1945:WSP

[279] Walter Rautenstrauch. What is scientific planning? Philosophy of Science, 12(1):8–18, January 1945. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184716.

Simms:1945:FMR

[280] Henry S. Simms. The future of medical research. Philosophy of Science, 12(1):19–22, January 1945. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184717.

Somerville:1945:SSD

[281] John Somerville. Soviet science and dialectical materialism. Philosophy of Science, 12(1):23–29, January 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184718.

Winthrop:1945:CDM

[282] Henry Winthrop. Conceptual difficulties in modern economic theory. Philosophy of Science, 12(1):30–39, January 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184719.

Anonymous:1945:FMa

[283] Anonymous. Front matter. Philosophy of Science, 12(1):??, January 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184714.

Lal:1945:PST

[284] Gobind Behari Lal. Popularization of science through news. Philosophy of Science, 12(2):41–44, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184276.

Reiser:1945:ISH

[285] Oliver L. Reiser. An institute of scientific humanism. Philosophy of Science, 12(2):45–51, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184277.

Feibleman:1945:NOE

[286] James Feibleman and Julius Friend. Normative organization and empirical fields. Philosophy of Science, 12(2):52–56, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184278.

Herrick:1945:NHE

[287] C. Judson Herrick. The natural history of experience. Philosophy of Science, 12(2):57–71, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184279.

Carnap:1945:IL

[288] Rudolf Carnap. On inductive logic. Philosophy of Science, 12(2):72– 97, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184280. See comment

[357].

Hempel:1945:DDC

[289] Carl G. Hempel and Paul Oppenheim. A definition of “degree of confirmation”. Philosophy of Science, 12(2):98–115, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184281.

Berenda:1945:BER

[290] Carlton W. Berenda. On Birkhoff’s and Einstein’s Relativity theory. Philosophy of Science, 12(2):116–119, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184282.

Hartung:1945:SFP

[291] Frank E. Hartung. The social function of positivism. Philosophy of Science, 12(2):120–133, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184283.

Zerby:1945:NDI

[292] Lewis Zerby. Normative, descriptive, and ideological elements in the writings of laski. Philosophy of Science, 12(2):134–145, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184284.

Carmichael:1945:FCN

[293] Peter A. Carmichael. Further concerning the null class. Philosophy of Science, 12(2):146, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184285.

Anonymous:1945:FMb

[294] Anonymous. Front matter. Philosophy of Science, 12(2):??, April 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184275.

Churchman:1945:PTB

[295] C. W. Churchman. Probability theory. I. Background. Philosophy of Science, 12(3):147–157, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184819.

Churchman:1945:PTIa

[296] C. W. Churchman. Probability theory. II. postulates of experimental method. Philosophy of Science, 12(3):158–164, July 1945. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184820.

Churchman:1945:PTIb

[297] C. W. Churchman. Probability theory. III. non-mechanical concepts. Philosophy of Science, 12(3):165–173, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184821.

Beck:1945:MFP

[298] Guido Beck. Mathematical formalism and the physical picture. Philosophy of Science, 12(3):174–178, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184822.

Witzemann:1945:CE

[299] Edgar J. Witzemann. Chemistry and evolution. Philosophy of Science, 12(3):179–189, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184823.

Kavanau:1945:TCF

[300] J. Lee Kavanau. A theory on causal factors in the origin of life. Philosophy of Science, 12(3):190–193, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184824.

Feibleman:1945:HEM

[301] James Feibleman. The hypothesis of esthetic measure. Philosophy of Science, 12(3):194–217, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184825.

Hull:1945:NCW

[302] Clark L. Hull. A note concerning Winthrop’s “conceptual difficulties in modern economic theory”. Philosophy of Science, 12(3):218, July 1945.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184826.

Churchman:1945:C

[303] C. W. Churchman and T. A. Cowan. A challenge. Philosophy of Science, 12(3):219–220, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184827.

Hawkins:1945:BRB

[304] David Hawkins. Book review: Theory of Games and Economic Behavior by John von Neumann; Oskar Morgenstern. Philosophy of Science, 12(3): 221–227, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184828.

Winthrop:1945:BRP

[305] Henry Winthrop. Book review: Philosophical analysis: The Philosophy of Bertrand Russell. Philosophy of Science, 12(3):227, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184829.

M:1945:BRBa

[306] W. M. M. Book review: The Appeal to Immediate Experience by Robert D. Mack. Philosophy of Science, 12(3):227–228, July 1945. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184830.

M:1945:BRBb

[307] W. M. M. Book review: The Science of Man in the World Crisis by Ralph Linton. Philosophy of Science, 12(3):228–229, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184831.

Anonymous:1945:FMc

[308] Anonymous. Front matter. Philosophy of Science, 12(3):??, July 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184818.

Malisoff:1945:PPS

[309] William Marias Malisoff. On the possible philosophies of science. Philosophy of Science, 12(4):231–236, October 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184249.

Johnson:1945:WTA

[310] A. H. Johnson. Whitehead’s theory of actual entities: Defence and criticism. Philosophy of Science, 12(4):237–295, October 1945. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184250.

Smith:1945:HVP

[311] Mapheus Smith. Hypothesis vs. problem in scientific investigation. Philosophy of Science, 12(4):296–301, October 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184251.

Bohnert:1945:SSC

[312] Herbert Gaylord Bohnert. The semiotic status of commands. Philosophy of Science, 12(4):302–315, October 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184252.

Rosenblueth:1945:RMS

[313] Arturo Rosenblueth and Norbert Wiener. The role of models in science. Philosophy of Science, 12(4):316–321, October 1945. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184253.

Anonymous:1945:VI

[314] Anonymous. Volume information. Philosophy of Science, 12(4):323–324, October 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184247.

Anonymous:1945:FMd

[315] Anonymous. Front matter. Philosophy of Science, 12(4):??, October 1945. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184248.

Malisoff:1946:NEA

[316] William Marias Malisoff. On the non-existence of the atomic secret. Philosophy of Science, 13(1):1–2, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184236.

Bentley:1946:LUP

[317] Arthur F. Bentley. Logicians’ underlying postulations. Philosophy of Science, 13(1):3–19, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184237.

Cowan:1946:MQ

[318] T. A. Cowan and C. W. Churchman. On the meaningfulness of questions. Philosophy of Science, 13(1):20–24, January 1946. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184238.

Storer:1946:LVI

[319] Thomas Storer. The logic of value imperatives. Philosophy of Science, 13(1):25–40, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184239.

Winn:1946:MN

[320] Ralph B. Winn. Mind and nature. Philosophy of Science, 13(1):41–52, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184240.

Thompson:1946:QHA

[321] Laura Thompson. In quest of an heuristic approach to the study of mankind. Philosophy of Science, 13(1):53–66, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184241.

Somerville:1946:OLB

[322] John Somerville. An open letter to Bertrand Russell. Philosophy of Science, 13(1):67–71, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184242.

Bergmann:1946:SCC

[323] Gustav Bergmann. Some comments on Carnap’s logic of induction. Philosophy of Science, 13(1):71–78, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184243.

Bahm:1946:OCP

[324] Archie J. Bahm. Opposition of compound propositions. Philosophy of Science, 13(1):78–79, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184244.

Feigl:1946:BRB

[325] Herbert Feigl. Book review: Ethics and Language by Charles L. Stevenson. Philosophy of Science, 13(1):80, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184245.

Anonymous:1946:BMa

[326] Anonymous. Back matter. Philosophy of Science, 13(1):??, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184246.

Anonymous:1946:FMa

[327] Anonymous. Front matter. Philosophy of Science, 13(1):??, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184235.

Anonymous:1946:VIa

[328] Anonymous. Volume information. Philosophy of Science, 13(1):??, January 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/184234.

Singer:1946:MVN

[329] Edgar A. Singer, Jr. Mechanism, vitalism, naturalism. A logico-historical study. Philosophy of Science, 13(2):81–99, April 1946. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184255.

McGill:1946:BAE

[330] V. J. McGill and Livingston Welch. A behaviorist analysis of emotions. Philosophy of Science, 13(2):100–122, April 1946. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184256.

Rusk:1946:SPP

[331] George Yeisley Rusk. Salvaging physiological psychology. Philosophy of Science, 13(2):123–130, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184257.

Ackoff:1946:TIC

[332] Russell L. Ackoff. Towards an interpretation of contemporary philosophy. Philosophy of Science, 13(2):131–136, April 1946. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184258.

Jones:1946:KEN

[333] Philip Chapin Jones. Kant, Euclid, and the non-Euclideans. Philosophy of Science, 13(2):137–143, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184259.

Miller:1946:WEC

[334] David L. Miller. Whitehead’s extensive continuum. Philosophy of Science, 13(2):144–149, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184260.

Phillips:1946:FE

[335] Dayton Phillips. The foundations of experience. Philosophy of Science, 13(2):150–165, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184261.

Malisoff:1946:SPP

[336] William Marias Malisoff. A science of the people, by the people and for the people. Philosophy of Science, 13(2):166–169, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184262.

Reid:1946:WD

[337] John R. Reid. What are definitions? Philosophy of Science, 13(2):170– 175, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184263.

Anonymous:1946:BMb

[338] Anonymous. Back matter. Philosophy of Science, 13(2):??, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184264.

Anonymous:1946:FMb

[339] Anonymous. Front matter. Philosophy of Science, 13(2):??, April 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184254.

Sellars:1946:PPR

[340] Roy Wood Sellars. The philosophy and physics of relativity. Philosophy of Science, 13(3):177–195, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184892.

vonBonin:1946:TSE

[341] Gerhardt von Bonin. Types and similitudes. An enquiry into the logic of comparative anatomy. Philosophy of Science, 13(3):196–202, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184893.

Eliasberg:1946:PP

[342] W. Eliasberg. Philosophy of psychotherapy. Philosophy of Science, 13(3): 203–214, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184894.

Randall:1946:CAS

[343] John A. Randall. Corism. Applied to specifying operations called scientific. Philosophy of Science, 13(3):215–222, July 1946. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184895.

Johnson:1946:WWW

[344] A. H. Johnson. The wit and wisdom of Whitehead. Philosophy of Science, 13(3):223–251, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184896.

Gentry:1946:EOP

[345] George Gentry. Eternal objects and the philosophy of organism. Philosophy of Science, 13(3):252–260, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184897.

Anonymous:1946:BMc

[346] Anonymous. Back matter. Philosophy of Science, 13(3):??, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184898.

Anonymous:1946:FMc

[347] Anonymous. Front matter. Philosophy of Science, 13(3):??, July 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184891.

Bergmann:1946:RR

[348] Gustav Bergmann. Remarks on realism. Philosophy of Science, 13(4):261– 273, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185206.

Mitchell:1946:MS

[349] E. T. Mitchell. Metaphysics and science. Philosophy of Science, 13(4): 274–280, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185207.

Miller:1946:MP

[350] David L. Miller. Metaphysics in physics. Philosophy of Science, 13(4):281– 286, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185208.

Churchman:1946:VU

[351] C. West Churchman and Russell L. Ackoff. Varieties of unification. Philosophy of Science, 13(4):287–300, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185209.

Burks:1946:PTA

[352] Arthur W. Burks. Peirce’s theory of abduction. Philosophy of Science, 13 (4):301–306, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185210.

Bowdery:1946:CFE

[353] George J. Bowdery. The concept of ‘field’ in electrical theory. Philosophy of Science, 13(4):307–324, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185211.

Black:1946:GJB

[354] Max Black. George J. Bowdery (1916–1945). Philosophy of Science, 13(4): 325, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185212.

Maslow:1946:PCV

[355] A. H. Maslow. Problem-centering vs. means-centering in science. Philosophy of Science, 13(4):326–331, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185213.

Meadows:1946:COA

[356] Paul Meadows. The cultural organization of action. Philosophy of Science, 13(4):332–338, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185214.

Churchman:1946:CBI

[357] C. West Churchman. Carnap’s “On Inductive Logic”. Philosophy of Science, 13(4):339–342, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185215. See [288].

Anonymous:1946:VIb

[358] Anonymous. Volume information. Philosophy of Science, 13(4):343–344, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185204.

Anonymous:1946:BMd

[359] Anonymous. Back matter. Philosophy of Science, 13(4):??, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185216.

Anonymous:1946:FMd

[360] Anonymous. Front matter. Philosophy of Science, 13(4):??, October 1946. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185205.

Miller:1947:BQM

[361] Cecil H. Miller. The basic question: Monism or dualism? Philosophy of Science, 14(1):1–12, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185529.

Berenda:1947:DPF

[362] Carlton W. Berenda. The determination of past by future events. A discussion of the Wheeler–Feynman absorption–radiation theory. Philosophy of Science, 14(1):13–19, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185530.

Klausner:1947:TDE

[363] Neal W. Klausner. Three decades of the epistemological dialectic 1900– 1930. Philosophy of Science, 14(1):20–43, January 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185531.

Witzemann:1947:SOL

[364] Edgar J. Witzemann. The scope, objectives and limitations of modern science as seen in the light of its history. Philosophy of Science, 14(1):44– 55, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185532.

Bawden:1947:PBD

[365] H. Heath Bawden. The psychical as a biological directive. Philosophy of Science, 14(1):56–67, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185533.

Hartung:1947:SFM

[366] Frank E. Hartung. Sociological foundations of modern science. Philosophy of Science, 14(1):68–95, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185534.

Mesthene:1947:NSE

[367] Emmanuel G. Mesthene. On the need for a scientific ethic. Philosophy of Science, 14(1):96–101, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185535.

M:1947:BRBa

[368] W. M. M. Book review: The Domain of Reality by William Gerber. Philosophy of Science, 14(1):102, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185536.

M:1947:BRBb

[369] W. M. M. Book review: Right Thinking by Edwin Arthur Burtt. Philosophy of Science, 14(1):102, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185537.

M:1947:BRBc

[370] W. M. M. Book review: The Idea of Nature by R. G. Collingwood. Philosophy of Science, 14(1):102–103, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185538.

M:1947:BRBd

[371] W. M. M. Book review: The Appeal to Immediate Experience by Robert D. Mack. Philosophy of Science, 14(1):103, January 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185539.

M:1947:BRBe

[372] W. M. M. Book review: Empirical Philosphies of Religion by James Alfred Martin. Philosophy of Science, 14(1):103, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185540.

M:1947:BRBf

[373] W. M. M. Book review: The a Priori in Physical Theory by Arthur Pap. Philosophy of Science, 14(1):103, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185541.

M:1947:BRBg

[374] W. M. M. Book review: The Authoritarian Attempt to Capture Education by John Dewey; Sidney Hook; Arthur E. Murphy; Irwin Edman; Bruce Bliven. Philosophy of Science, 14(1):103–104, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185542.

Anonymous:1947:FMa

[375] Anonymous. Front matter. Philosophy of Science, 14(1):??, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185528.

Anonymous:1947:VIa

[376] Anonymous. Volume information. Philosophy of Science, 14(1):??, January 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185527.

Biser:1947:EAP

[377] Erwin Biser. Entity and aspects. (as pertaining to physical theory). Philosophy of Science, 14(2):105–115, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185834.

Smith:1947:NPG

[378] Mapheus Smith. A note on the progressive generalization of data. Philosophy of Science, 14(2):116–122, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185835.

Hall:1947:NPV

[379] Everett W. Hall. On the nature of the predicate, ’verified’. Philosophy of Science, 14(2):123–131, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185836.

King:1947:WCM

[380] Hugh R. King. A. N. Whitehead and the concept of metaphysics. Philosophy of Science, 14(2):132–151, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185837.

Bergmann:1947:SDL

[381] Gustav Bergmann. Sense data, linguistic conventions, and existence. Philosophy of Science, 14(2):152–163, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185838.

Jones:1947:CP

[382] Philip Chapin Jones. Communication in philosophy. Philosophy of Science, 14(2):164–170, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185839.

M:1947:BRBh

[383] W. M. M. Book review: Modern Man Is Obsolete by Norman Cousins. Philosophy of Science, 14(2):171, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185840.

M:1947:BRBi

[384] W. M. M. Book review: Science and Freedom by Lyman Bryson. Philosophy of Science, 14(2):171, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185841.

M:1947:BRBj

[385] W. M. M. Book review: Science and the Planned State by John R. Baker. Philosophy of Science, 14(2):171–172, April 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185842.

M:1947:BRBk

[386] W. M. M. Book review: Soviet Philosophy by John Somerville. Philosophy of Science, 14(2):172, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185843.

M:1947:BRBl

[387] W. M. M. Book review: Scientific Man vs. Power Politics by Hans J. Morgenthau. Philosophy of Science, 14(2):172–173, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185844.

M:1947:BRBm

[388] W. M. M. Book review: The Source of Human Good by Henry N. Wieman. Philosophy of Science, 14(2):173, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185845.

M:1947:BRBn

[389] W. M. M. Book review: From Max Weber; Essays in Sociology by H. H. Gerth; C. W. Mills. Philosophy of Science, 14(2):173, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185846.

M:1947:BRBo

[390] W. M. M. Book review: The Geometry of Art and Life by Matila Ghyka. Philosophy of Science, 14(2):173–174, April 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185847.

M:1947:BRBp

[391] W. M. M. Book review: People in Quandaries by Wendell Johnson. Philosophy of Science, 14(2):174, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185848.

M:1947:BRBq

[392] W. M. M. Book review: Wisdom with Power by Aaron Hyam. Philosophy of Science, 14(2):174, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185849.

M:1947:BRBr

[393] W. M. M. Book review: What Is Life? by Erwin Schroedinger. Philosophy of Science, 14(2):174, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185850.

M:1947:BRBs

[394] W. M. M. Book review: Time, Knowledge and the Nebulae by Martin Johnson. Philosophy of Science, 14(2):175, April 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185851.

M:1947:BRBt

[395] W. M. M. Book review: The University at the Crossroads by Henry E. Sigerist. Philosophy of Science, 14(2):175, April 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185852.

Churchman:1947:MAA

[396] C. West Churchman. Much ado about probability. Philosophy of Science, 14(2):176–178, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185853.

Anonymous:1947:FMb

[397] Anonymous. Front matter. Philosophy of Science, 14(2):??, April 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185833.

Fries:1947:SCV

[398] Horace S. Fries. Science, causation, and value. Philosophy of Science, 14(3):179–180, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185914.

Sellars:1947:PPE

[399] Wilfrid Sellars. Pure pragmatics and epistemology. Philosophy of Science, 14(3):181–202, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185915.

Demos:1947:DAE

[400] Raphael Demos. Doubts about empiricism. Philosophy of Science, 14(3): 203–218, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185916.

Miller:1947:NSS

[401] David L. Miller. The nature of scientific statements. Philosophy of Science, 14(3):219–223, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185917.

Williams:1947:SK

[402] Elgin Williams. Sociologists and knowledge. Philosophy of Science, 14(3): 224–230, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185918.

Klee:1947:LAS

[403] James B. Klee. Learning — acquisition or selection — possibility versus probability. Philosophy of Science, 14(3):231–241, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185919.

Moulyn:1947:MMP

[404] Adrian C. Moulyn. Mechanisms and mental phenomena. Philosophy of Science, 14(3):242–253, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185920.

Herrmann:1947:MFA

[405] Heinz Herrmann. Morphological and functional aspects of living matter and Whitehead’s category of actual entity. Philosophy of Science, 14(3): 254–260, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185921.

Lapan:1947:ICB

[406] Arthur Lapan. Incompatibilities and conflicts: Breakdown. Philosophy of Science, 14(3):261–265, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185922.

Lillie:1947:BDP

[407] Ralph S. Lillie. Biological directiveness and the psychical. A note. Philosophy of Science, 14(3):266–268, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185923.

Churchman:1947:ES

[408] C. West Churchman and Russell L. Ackoff. Ethics and science. Philosophy of Science, 14(3):269–271, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185924.

Bergmann:1947:PPP

[409] Gustav Bergmann. Philosophical and psychological pragmatics. Philosophy of Science, 14(3):271–273, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185925.

Anonymous:1947:FMc

[410] Anonymous. Front matter. Philosophy of Science, 14(3):??, July 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185913.

Biser:1947:MRP

[411] Erwin Biser and Enos E. Witmer. Methodology of research and progress in science. Philosophy of Science, 14(4):275–288, October 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184987.

White:1947:LMR

[412] Leslie A. White. The locus of mathematical reality: An anthropological footnote. Philosophy of Science, 14(4):289–303, October 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184988.

Churchman:1947:EMP

[413] C. West Churchman and Russell L. Ackoff. An experimental measure of personality. Philosophy of Science, 14(4):304–332, October 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184989.

Hall:1947:CAV

[414] Everett W. Hall. A categorial analysis of value. Philosophy of Science, 14 (4):333–344, October 1947. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184990.

Somerville:1947:ESS

[415] John Somerville. Ethics and social science: Case history of a sharp practice. Philosophy of Science, 14(4):345–347, October 1947. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184991.

Anonymous:1947:VIb

[416] Anonymous. Volume information. Philosophy of Science, 14(4):349–350, October 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184985.

Anonymous:1947:FMd

[417] Anonymous. Front matter. Philosophy of Science, 14(4):??, October 1947. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184986.

Frank:1948:MDW

[418] Philip Frank and C. West Churchman. In memoriam: Dr. William M. Malisoff. Philosophy of Science, 15(1):1–3, January 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185193.

Hinshaw:1948:ERS

[419] Virgil G. Hinshaw, Jr. Epistemological Relativism and the sociology of knowledge. Philosophy of Science, 15(1):4–10, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185194.

Altschul:1948:VUM

[420] Eugen Altschul and Erwin Biser. The validity of unique mathematical models in science. Philosophy of Science, 15(1):11–24, January 1948. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185195.

Schlegel:1948:APP

[421] Richard Schlegel. Atemporal processes in physics. Philosophy of Science, 15(1):25–35, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185196.

Feibleman:1948:SPD

[422] James Feibleman. A set of postulates and a definition for science. Philosophy of Science, 15(1):36–38, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185197.

Leverett:1948:MWM

[423] Hollis M. Leverett. The mind we measure and its dimensions. Philosophy of Science, 15(1):39–46, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185198.

Bell:1948:MRP

[424] Charles G. Bell. Mechanistic replacement of purpose in biology. Philosophy of Science, 15(1):47–51, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185199.

Miller:1948:ECE

[425] David L. Miller. The effect of the concept of evolution on scientific methodology. Philosophy of Science, 15(1):52–60, January 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185200.

Klee:1948:PD

[426] James B. Klee. A point of departure. Philosophy of Science, 15(1):61–70, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185201.

Feuer:1948:MIS

[427] Lewis S. Feuer. Materialism, idealism and science. Philosophy of Science, 15(1):71–75, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185202.

Gregory:1948:PSS

[428] Ralph Gregory. The “psychical” as secondary and as secret. Philosophy of Science, 15(1):76–79, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185203.

Anonymous:1948:FMa

[429] Anonymous. Front matter. Philosophy of Science, 15(1):??, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185192.

Anonymous:1948:VIa

[430] Anonymous. Volume information. Philosophy of Science, 15(1):??, January 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185191.

Anonymous:1948:E

[431] Anonymous. Editorial. Philosophy of Science, 15(2):81–82, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185158.

Adams:1948:IR

[432] John S. Adams, Jr. Introductory remarks. Philosophy of Science, 15(2): 83–84, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185159.

Olmstead:1948:STB

[433] Paul S. Olmstead. Some thoughts on “What the Natural Scientist Needs from the Social Scientist”. Philosophy of Science, 15(2):85–86, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185160.

Phelps:1948:WPS

[434] Everett R. Phelps. What the physical scientists need from the social scientists. Philosophy of Science, 15(2):87–89, April 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185161.

Orens:1948:PSS

[435] Irving P. Orens. Physical science and the social sciences. Philosophy of Science, 15(2):90–95, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185162.

Preston:1948:CEC

[436] Malcolm G. Preston. Concerning an essential condition of cooperative work. Philosophy of Science, 15(2):96–99, April 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185163.

Butts:1948:SSR

[437] Carl F. Butts. Science and social responsibility. Philosophy of Science, 15(2):100–103, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185164.

Sellars:1948:DNS

[438] Roy Wood Sellars. Do the natural sciences have need of the social sciences? Philosophy of Science, 15(2):104–108, April 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185165.

Hartung:1948:CSP

[439] Frank E. Hartung. On the contribution of sociology to the physical sciences. Philosophy of Science, 15(2):109–115, April 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185166.

Ackoff:1948:D

[440] Russell L. Ackoff. Discussion. Philosophy of Science, 15(2):116–117, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185167.

Lillie:1948:SAT

[441] Ralph S. Lillie. Some aspects of theoretical biology. Philosophy of Science, 15(2):118–134, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185168.

Hempel:1948:SLE

[442] Carl G. Hempel and Paul Oppenheim. Studies in the logic of explanation. Philosophy of Science, 15(2):135–175, April 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185169. See comments [463].

Anonymous:1948:NNa

[443] Anonymous. News and notices. Philosophy of Science, 15(2):176–177, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185170.

Anonymous:1948:FMb

[444] Anonymous. Front matter. Philosophy of Science, 15(2):??, April 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185157.

Agar:1948:WLO

[445] W. E. Agar. The wholeness of the living organism. Philosophy of Science, 15(3):179–191, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185082.

Wolff:1948:UGT

[446] Kurt H. Wolff. The unique and the general: Toward a philosophy of sociology. Philosophy of Science, 15(3):192–210, July 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185083.

Williams:1948:ATS

[447] Gardner Williams. Absolute truth and the shadow of doubt. Philosophy of Science, 15(3):211–224, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185084.

Feuer:1948:IED

[448] Lewis S. Feuer. Indeterminacy and economic development. Philosophy of Science, 15(3):225–241, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185085.

Becker:1948:SPA

[449] Arthur Peter Becker. Some philosophical aspects of economics. Philosophy of Science, 15(3):242–246, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185086.

Ledgerwood:1948:TNG

[450] Richard Ledgerwood. Taxonomic note on the gnosology of modern science. Philosophy of Science, 15(3):247–248, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185087.

Churchman:1948:SPI

[451] C. West Churchman. Statistics, pragmatics, induction. Philosophy of Science, 15(3):249–268, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185088. See note [494] and reply [495].

Strausz-Hupe:1948:BRB

[452] Robert Strausz-Hupé. Book review: The Open Society and Its Enemies by K. R. Popper. Philosophy of Science, 15(3):269–271, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185089.

Ackoff:1948:BRBa

[453] Russell L. Ackoff. Book review: The Logic of the Sciences and the Humanities by F. S. C. Northrop. Philosophy of Science, 15(3):271–272, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185090.

Ackoff:1948:BRBb

[454] Russell L. Ackoff. Book review: Lenin as Philosopher by Anton Pannekoek. Philosophy of Science, 15(3):272–273, July 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185091.

Anonymous:1948:NNb

[455] Anonymous. News and notices. Philosophy of Science, 15(3):274, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185092.

Anonymous:1948:FMc

[456] Anonymous. Front matter. Philosophy of Science, 15(3):??, July 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185081.

Frank:1948:PLM

[457] Philipp Frank. The place of logic and metaphysics in the advancement of modern science. Philosophy of Science, 15(4):275–286, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184961.

Sellars:1948:CIL

[458] Wilfrid Sellars. Concepts as involving laws and inconceivable without them. Philosophy of Science, 15(4):287–315, October 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184962.

Storer:1948:PRB

[459] Thomas Storer. The philosophical relevance of a “behavioristic semiotic”. Philosophy of Science, 15(4):316–330, October 1948. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184963. See comments [460].

Morris:1948:CMS

[460] Charles Morris. Comments on Mr. Storer’s paper. Philosophy of Science, 15(4):330–332, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184964.

See [459].

Williams:1948:CWS

[461] Elgin Williams. Can we save science? Philosophy of Science, 15(4):333– 341, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184965.

Lundberg:1948:RCW

[462] George A. Lundberg. Rejoinder to “can we save science?”. Philosophy of Science, 15(4):341–347, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184966.

Miller:1948:CBS

[463] David L. Miller. Comments on “Studies in the Logic of Explanation”. Philosophy of Science, 15(4):348–349, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184967. See [442].

Hempel:1948:RDM

[464] Carl G. Hempel and Paul Oppenheim. Reply to David L. Miller’s comments. Philosophy of Science, 15(4):350–352, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184968.

Bergmann:1948:DNC

[465] Gustav Bergmann. Descriptions in nonextensional contexts. Philosophy of Science, 15(4):353–355, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184969.

Benjamin:1948:BRB

[466] A. Cornelius Benjamin. Book review: Logic and Scientific Method by Herbert L. Searles. Philosophy of Science, 15(4):356, October 1948. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184970.

Anonymous:1948:BM

[467] Anonymous. Back matter. Philosophy of Science, 15(4):357–361, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184972.

Anonymous:1948:NNc

[468] Anonymous. News and notices. Philosophy of Science, 15(4):361, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184971.

Anonymous:1948:VIb

[469] Anonymous. Volume information. Philosophy of Science, 15(4):363–364, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184959.

Anonymous:1948:FMd

[470] Anonymous. Front matter. Philosophy of Science, 15(4):??, October 1948. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184960.

Anonymous:1949:EPS

[471] Anonymous. Editorial: Philosophy of science and liberalism. Philosophy of Science, 16(1):1–2, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185685.

Frank:1949:PED

[472] Jerome Frank. The place of the expert in a democratic society. Philosophy of Science, 16(1):3–24, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185686.

Burrow:1949:SNS

[473] Trigant Burrow. The social neurosis: A study in “clinical anthropology”. Philosophy of Science, 16(1):25–40, January 1949. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185687.

Rudner:1949:FNF

[474] Richard Rudner. Formal and non-formal. Philosophy of Science, 16(1):41– 48, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185688.

Jones:1949:SP

[475] Philip Chapin Jones. Subjectivity in philosophy. Philosophy of Science, 16(1):49–57, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185689.

Greenberg:1949:LAK

[476] Joseph H. Greenberg. The logical analysis of kinship. Philosophy of Science, 16(1):58–64, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185690.

Copilowish:1949:LAM

[477] Irving M. Copilowish. Language analysis and metaphysical inquiry. Philosophy of Science, 16(1):65–70, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185691.

Bergmann:1949:TCI

[478] Gustav Bergmann. Two criteria for an ideal language. Philosophy of Science, 16(1):71–74, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185692.

Copilowish:1949:RPB

[479] Irving M. Copilowish. Reply to Professor Bergmann. Philosophy of Science, 16(1):74, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185693.

Brodbeck:1949:CTR

[480] May Brodbeck. Coherence theory reconsidered: Professor Werkmeister on semantics and on the nature of empirical laws. Philosophy of Science, 16(1):75–85, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185694.

Hoekstra:1949:BRB

[481] Raymond Hoekstra. Book review: The Philosophy of Francis Bacon by Fulton H. Anderson. Philosophy of Science, 16(1):86–87, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185695.

Ackoff:1949:BRBa

[482] R. L. Ackoff. Book review: Projective Methods by Lawrence K. Frank. Philosophy of Science, 16(1):87, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185696.

Churchman:1949:BRBa

[483] C. West Churchman. Book reviews: Techniques of Statistical Analysis by Statistical Research Group; Sampling Inspection by Statistical Research Group. Philosophy of Science, 16(1):88, January 1949. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185697.

Churchman:1949:BRBb

[484] C. West Churchman. Book review: The Mathematical Analysis of Logic by George Boole. Philosophy of Science, 16(1):88, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185698.

Anonymous:1949:FMa

[485] Anonymous. Front matter. Philosophy of Science, 16(1):??, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185684.

Anonymous:1949:VIa

[486] Anonymous. Volume information. Philosophy of Science, 16(1):??, January 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185683.

Ruddick:1949:HSL

[487] Chester T. Ruddick. Hume on scientific law. Philosophy of Science, 16(2): 89–93, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185127.

Henle:1949:FFS

[488] Paul Henle. On the fourth figure of the syllogism. Philosophy of Science, 16(2):94–104, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185128.

Feuer:1949:DMS

[489] Lewis S. Feuer. Dialectical materialism and Soviet science. Philosophy of Science, 16(2):105–124, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185129.

Metzger:1949:IIS

[490] Walter P. Metzger. Ideology and the intellectual: A study of thorstein Veblen. Philosophy of Science, 16(2):125–133, April 1949. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185130.

Ballard:1949:PM

[491] Edward G. Ballard. The paradox of measurement. Philosophy of Science, 16(2):134–136, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185131.

Fitch:1949:PMS

[492] Frederic B. Fitch. The problem of the morning star and the evening star. Philosophy of Science, 16(2):137–141, April 1949. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185132.

Freund:1949:SVP

[493] John E. Freund. Statistical vs. pragmatic inference. Philosophy of Science, 16(2):142–147, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185133.

Cowan:1949:NCB

[494] Thomas A. Cowan. A note on Churchman’s “Statistics, Pragmatics, Induction”. Philosophy of Science, 16(2):148–150, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185134. See [451, 495].

Churchman:1949:RCB

[495] C. West Churchman. Reply to comments on “Statistics, Pragmatics, Induction”. Philosophy of Science, 16(2):151–153, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185135. See [451, 494].

Ackoff:1949:EPP

[496] Russell L. Ackoff. An educational program for the philosophy of science. Philosophy of Science, 16(2):154–157, April 1949. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185136.

Ruddick:1949:BRB

[497] C. T. Ruddick. Book review: The Solar System Analysed by F. C. Attwood. Philosophy of Science, 16(2):158, April 1949. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185137.

Benjamin:1949:BRB

[498] A. Cornelius Benjamin. Book review: Reflections on the Philosophy of Sir Arthur Eddington by A. D. Ritchie. Philosophy of Science, 16(2):158– 159, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185138.

Ackoff:1949:BRBb

[499] Russell L. Ackoff. Book review: Cybernetics by Norbert Wiener. Philosophy of Science, 16(2):159–160, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185139.

Anonymous:1949:FMb

[500] Anonymous. Front matter. Philosophy of Science, 16(2):??, April 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185126.

Merton:1949:RAS

[501] Robert K. Merton. The role of applied social science in the formation of policy: A research memorandum. Philosophy of Science, 16(3):161– 181, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185512.

Bain:1949:NSV

[502] Read Bain. Natural science and value-policy. Philosophy of Science, 16 (3):182–192, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185513.

Barkin:1949:ASS

[503] Solomon Barkin. Applied social science in the American trade-union movement. Philosophy of Science, 16(3):193–197, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185514.

Cartwright:1949:BAS

[504] Dorwin Cartwright. Basic and applied social psychology. Philosophy of Science, 16(3):198–208, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185515.

Hauser:1949:SSS

[505] Philip M. Hauser. Social science and social engineering. Philosophy of Science, 16(3):209–218, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185516.

Shils:1949:SSS

[506] E. A. Shils. Social science and social policy. Philosophy of Science, 16(3): 219–242, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185517.

Ulrich:1949:CMA

[507] David N. Ulrich. A clinical method in applied social science. Philosophy of Science, 16(3):243–249, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185518.

Rusk:1949:GMG

[508] George Y. Rusk. General mensurational gestaltism. Philosophy of Science, 16(3):250–259, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185519.

Wiener:1949:SCD

[509] Norbert Wiener. Sound communication with the deaf. Philosophy of Science, 16(3):260–262, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185520.

Hallowell:1949:BRB

[510] John H. Hallowell. Book review: The Liberal Spirit. Essays on Problems of Freedom in the Modern World by Horace M. Kallen. Philosophy of Science, 16(3):263–264, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185521.

Ackley:1949:BRB

[511] Sheldon Ackley. Book review: Integration of the Humanities and the Social Sciences: A Symposium by Ernest E. Leisy. Philosophy of Science, 16(3): 264–265, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185522.

Biser:1949:BRB

[512] E. Biser. Book review: Power and Events by A. P. Ushenko. Philosophy of Science, 16(3):265–266, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185523.

Ackoff:1949:BRBc

[513] Russell L. Ackoff. Book review: The Scientific Attitude by C. H. Waddington. Philosophy of Science, 16(3):266, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185524.

Ackoff:1949:BRBd

[514] Russell L. Ackoff. Book review: Readings in Philosophical Analysis by Herbert Feigl; Wilfrid Sellars. Philosophy of Science, 16(3):266–267, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185525.

Anonymous:1949:NNa

[515] Anonymous. News and notices. Philosophy of Science, 16(3):268, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185526.

Anonymous:1949:FMc

[516] Anonymous. Front matter. Philosophy of Science, 16(3):??, July 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185511.

Jordan:1949:PMQ

[517] P. Jordan. On the process of measurement in quantum mechanics. Philosophy of Science, 16(4):269–278, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185067.

Lenzen:1949:CRQ

[518] Victor F. Lenzen. Concepts and reality in quantum mechanics. Philosophy of Science, 16(4):279–286, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185068.

Margenau:1949:RQM

[519] Henry Margenau. Reality in quantum mechanics. Philosophy of Science, 16(4):287–302, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185069.

Menzel:1949:PPQ

[520] Donald H. Menzel and David Layzer. The physical principles of the quantum theory. Philosophy of Science, 16(4):303–324, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185070.

Reiser:1949:REW

[521] Oliver L. Reiser. A resolution of the “East-West Problem” by way of a scientific humanism. Philosophy of Science, 16(4):325–335, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185071.

Geiger:1949:NNF

[522] George R. Geiger. A note on the naturalistic fallacy. Philosophy of Science, 16(4):336–342, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185072.

Linsky:1949:SNC

[523] Leonard Linsky. Some notes on Carnap’s concept of intensional isomorphism and the paradox of analysis. Philosophy of Science, 16(4):343–347, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185073.

Carnap:1949:RLL

[524] Rudolf Carnap. A reply to Leonard Linsky. Philosophy of Science, 16 (4):347–350, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185074.

Sheppard:1949:BRB

[525] Harold L. Sheppard. Book review: Scientists and Amateurs: A History of the Royal Society by Dorothy Stimson. Philosophy of Science, 16(4):351, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185075.

A:1949:BRB

[526] R. L. A. Book review: Psychology and Ethics by Harry L. Hollingworth. Philosophy of Science, 16(4):351–352, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185076.

Minas:1949:BRB

[527] J. Sayer Minas. Book review: Christian Science and Philosophy by Henry W. Steiger. Philosophy of Science, 16(4):352–353, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185077.

H:1949:BRBa

[528] F. H. Book review: Science, Faith and Society by Michael Polanyi. Philosophy of Science, 16(4):353, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185078.

Anonymous:1949:NNb

[529] Anonymous. News and notices. Philosophy of Science, 16(4):354, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185080.

H:1949:BRBb

[530] F. H. Book review: Science since 1500 by H. T. Pledge. Philosophy of Science, 16(4):354, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185079.

Anonymous:1949:VIb

[531] Anonymous. Volume information. Philosophy of Science, 16(4):355–356, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185065.

Anonymous:1949:FMd

[532] Anonymous. Front matter. Philosophy of Science, 16(4):??, October 1949. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185066.